### v1

# A subcopula characterization of dependence for the Multivariate Bernoulli Distribution

Abstract
By applying Sklar’s theorem to the Multivariate Bernoulli Distribution (MBD), it is proposed a framework that decouples the marginal distributions from the dependence structure, providing a clearer understanding of how binary variables interact. Explicit formulas are derived under the MBD using subcopulas to introduce dependence measures for interactions of all orders, not just pairwise. A bayesian inference approach is also applied to estimate the parameters of the MBD, offering practical tools for parameter estimation and dependence analysis in real-world applications. The results obtained contribute to the application of subcopulas of multivariate binary data, with a real data example of comorbidities in COVID-19 patients.
Keywords: Sklar theorem, subcopula, dependence, Bernoulli, binary data.
1 Introduction
In statistical modeling, the analysis of multivariate binary data often requires flexible methods to capture the dependence structure between variables. The Multivariate Bernoulli Distribution (MBD) offers a natural framework for modeling such data, yet specifying and understanding the dependence between variables in this context can be challenging. This is especially true when working with multivariate data where traditional correlation measures fall short due to the binary nature of the variables and the dimensionality.
A representation of the MBD that provides an alternative to the traditional log-linear model for binary variables was proposed by [17] using the concept of Kronecker product from matrix calculus, in terms of $12^{n}-$ 1 parameters, where $n$ is the dimension of a random vector with MBD, but where the dependency vector of parameters are, in fact, non standardized central moments. In [11] several flexible methods for simulating random binary sequences with fixed marginal distributions and specified degrees of association between the variables are discussed, but only pairwise dependencies are considered. An algorithm is derived in [7] for generating systems of correlated binary data, but allowing for the specification of just pairwise correlations within each system. A simple method to characterize multivariate Bernoulli variables with given means is investigated in [5] but only pairwise are correlations are discussed.
Copulas, and more generally subcopulas, provide a straightforward approach to describing the dependence structure in multivariate settings by decoupling the marginal distributions from the joint distribution. Introduced by Sklar’s theorem [15], subcopulas extend copula theory to cases where the marginals are not necessarily continuous, making them particularly useful for multivariate discrete data. This paper focuses on leveraging subcopulas to characterize the dependence within the MBD, offering a framework to better understand and quantify dependence of all orders (pariwise, three-wise, four-wise, etc.)
Building upon previous work in copula theory and dependence measures, the present work introduces a formal subcopula-based characterization of the dependence structure within the MBD, and illustrates its usefulness to tackle the compatibility problem of building multivariate models in terms of lower dimension marginals.
This work provides a systematic approach for analyzing the joint dependence of multivariate binary data, with potential applications in fields ranging from genetics to econometrics where binary outcomes are prevalent. The results presented here not only contribute to the theoretical understanding of subcopulas but also offer practical tools for applied statistics and data analysis, and Julia programming [2] code is provided for their computational implementation and application.
2 Subcopulas and dependence
Sklar’s Theorem [15] proves that for any $n-$dimensional vector of random variables $(X_{1},\ldots,X_{n})$ there exists a functional link $S^{(n)}$ between its probability joint distribution $F_{X_{1},\ldots,X_{n}}(x_{1},\ldots,x_{n})=\mathbb{P}(X_{1}\leq x_{1},\ldots,X_{n}\leq x_{n})$ and its corresponding univariate marginal distributions $F_{X_{i}}(x_{i})=\mathbb{P}(X_{i}\leq x_{i}),$ $i\in\{1,\ldots,n\}:$
$F_{X_{1},\ldots,X_{n}}(x_{1},\ldots,x_{n})\,=\,S^{(n)}\left(F_{X_{1}}(x_{1}),\ldots,F_{X_{n}}(x_{n})\right)\,,$ (1)
for any $(x_{1},\ldots,x_{n})\in\overline{\mathbb{R}}^{n},$ and where the function $S^{(n)}:D_{1}\times\cdots\times D_{n}\rightarrow[0,1]$ is unique, with $D_{i}=\mbox{Ran}\,F_{X_{i}}$ for $i\in\{1,\ldots,n\},$ which implies $\{0,1\}\subseteq D_{i}\subseteq[0,1].$ Such link $S^{(n)}$ is called a subcopula function, and in the particular case where every $D_{i}=[0,1]$ then it is called a copula function, which would be the case when all the random variables are continuous. If at least one of the random variables is not continuous then the domain of the underlying subcopula $S^{(n)}$ is a proper subset of the unit hypercube $[0,1]^{n}.$ A formal definition for subcopula functions and their properties are discussed in detail in [13] and [12].
Recalling that the random variables $X_{1},\ldots,X_{n}$ are independent if and only if their probability joint distribution is equal to the product of its univariate marginals, then as an immediate consequence of (1) the unique subcopula function under independence is:
$\Pi^{(n)}(u_{1},\ldots,u_{n}):=u_{1}\cdots u_{n}\,,\qquad(u_{1},\ldots,u_{n})\in D_{1}\times\cdots\times D_{n}\,.$ (2)
Moreover, as a consequence of the Fréchet-Hoeffding bounds for joint distributions (see [6] and [8]) and applying (1) we may obtain bounds for any subcopula $S^{(n)}:$
$W^{(n)}(u_{1},\ldots,u_{n})\,\leq\,S^{(n)}(u_{1},\ldots,u_{n})\,\leq\,M^{(n)}(u_{1},\ldots,u_{n})\,,$ (3)
where $W^{(n)}(u_{1},\ldots,u_{n}):=\max\{u_{1}+\cdots+u_{n}-n+1,0\}$ and $M^{(n)}(u_{1},\ldots,u_{n}):=\min\{u_{1},\ldots,u_{n}\}.$ The upper bound $M^{(n)}$ is always a subcopula, but the lower bound $W^{(n)}$ is subcopula only for $n=2,$ though it is still the best possible lower bound for $n>2,$ see [12].
From any $n-$dimensional subcopula $S^{(n)}(u_{1},\ldots,u_{n})$ and $2\leq k<n$ it is possible to obtain $\binom{n}{k}$ $k-$dimensional marginal subcopulas $S^{(k)}(v_{1},\ldots,v_{k})$ defined in terms of $S^{(n)}(u_{1},\ldots,u_{n})$ where $k$ of its entries are equal to exactly one of each $\{v_{1},\ldots,v_{k}\}$ and the remaining $n-k$ are equal to $1$
Since the univariate marginal distributions $F_{X_{i}}(x_{i})=\mathbb{P}(X_{i}\leq x_{i})$ have no information about how each random variable $X_{i}$ interacts with others, then as a consequence of (1) all the information about the dependence among the random variables $X_{1},\ldots,X_{n}$ is contained in their unique underlying subcopula $S^{(n)},$ and therefore any attempt to measure degrees or intensity of dependence should extract information from $S^{(n)}.$ For the particular case of bivariate copulas (when both random variables are continuous) several copula-based measures have been studied, such as the concordance measures by Spearman [16] or Kendall [10], or dependence measures by Schweizer-Wolff [14] or Hoeffding [8]. For the case of proper subcopulas (that are not copulas) a subcopula dependence measure has been proposed by Erdely [3].
3 Bivariate Bernoulli
[…]
Accordingly to (1) the domain of the underlying subcopula in this case is the set $\mbox{Dom}\,S^{(2)}=\{0,\theta_{1},1\}\times\{0,\theta_{2},1\}.$ From the formal definition of bivariate subcopula (see [12], for example) it always satisfies $S^{(2)}(u,0)=0=S^{(2)}(0,v),$ $S^{(2)}(u,1)=u,$ and $S^{(2)}(1,v)=v,$ and therefore:
$S^{(2)}(u,v)=\begin{cases}\,1&\mbox{ if }(u,v)=(1,1),\\ \,\theta_{1}&\mbox{ if }(u,v)=(\theta_{1},1),\\ \,\theta_{2}&\mbox{ if }(u,v)=(1,\theta_{2}),\\ \,\theta_{12}&\mbox{ if }(u,v)=(\theta_{1},\theta_{2}),\\ \,0&\mbox{ elsewhere,}\end{cases}$ (4)
where $\theta_{12}$ is any value satisfying:
$\max\{\theta_{1}+\theta_{2}-1,0\}\,\leq\,\theta_{12}\,\leq\,\min\{\theta_{1},\theta_{2}\}$ (5)
as a consequence of (3), that in this case acts as a bivariate dependence parameter. Since $\mbox{Ran}\,(X,Y)=\{(0,0),(0,1),(1,0),(1,1)\}$ let $\{p_{00},p_{01},p_{10},p_{11}\}$ be the joint point probabilities:
$p_{ij}:=\mathbb{P}(X_{1}=i,X_{2}=j)\,,\qquad i,j\in\{0,1\},$ (6)
where necessarily $0\leq p_{ij}\leq 1$ and $p_{00}+p_{01}+p_{10}+p_{11}=1,$ and therefore only three of the $p_{ij}$ values need to be specified, since the fourth is just $1$ minus the sum of the other three. Applying (1) and (4) to calculate each $p_{ij}$ in terms of the joint distribution function $F_{12}(i,j)=\mathbb{P}(X_{1}\leq i,X_{2}\leq j):$
$p_{ij}$ $=$ $F_{12}(i,j)-F_{12}(i,j-1)-F_{12}(i-1,j)+F_{12}(i-1,j-1)\,,$ (7)
$=$ $S^{(2)}\left(F_{1}(i),F_{2}(j)\right)-S^{(2)}\left(F_{1}(i),F_{2}(j-1)\right)-S^{(2)}\left(F_{1}(i-1),F_{2}(j)\right)\ldots$
$+S^{(2)}\left(F_{1}(i-1),F_{2}(j-1)\right)\,,$
and therefore:
$p_{00}$ $=$ $F_{12}(0,0)=S^{(2)}(\theta_{1},\theta_{2})=\theta_{12}\,,$ (8)
$p_{01}$ $=$ $F_{12}(0,1)-F_{12}(0,0)=S^{(2)}(\theta_{1},1)-S^{(2)}(\theta_{1},\theta_{2})=\theta_{1}-\theta_{12}\,,$
$p_{10}$ $=$ $F_{12}(1,0)-F_{12}(0,0)=S^{(2)}(1,\theta_{2})-S^{(2)}(\theta_{1},\theta_{2})=\theta_{2}-\theta_{12}\,,$
$p_{11}$ $=$ $1-\theta_{1}-\theta_{2}+\theta_{12}\,.$
Formulas (8) allow to fully characterize the family $\mathcal{B}_{2}(\theta_{1},\theta_{2},\theta_{12})$ of all bivariate Bernoulli distributions in terms of the marginal parameters $10<\theta_{1}<$ 1 and $0<\theta_{2}<1,$ and a bivariate dependence parameter $\theta_{12}$ that must satisfy $\max\{\theta_{1}+\theta_{2}-1,0\}\leq\theta_{12}\leq\min\{\theta_{1},\theta_{2}\}.$ The subcopula based dependence measure proposed by [3] for $\mathcal{B}_{2}(\theta_{1},\theta_{2},\theta_{12})$ becomes:
$\mu(X_{1},X_{2})=\begin{cases}\frac{\theta_{12}-\theta_{1}\theta_{2}}{\min\{\theta_{1},\theta_{2}\}-\theta_{1}\theta_{2}}&\mbox{ if }\theta_{12}\geq\theta_{1}\theta_{2}\,,\\ {}&{}\textstyle\\ \frac{\theta_{12}-\theta_{1}\theta_{2}}{\theta_{1}\theta_{2}-\max\{\theta_{1}+\theta_{2}-1,0\}}&\mbox{ if }\theta_{12}<\theta_{1}\theta_{2}\,.\end{cases}$ (9)
Two Bernoulli random variables are independent if and only if $\theta_{12}=\theta_{1}\theta_{2}$ as an immediate consequence of (2), or equivalently if and only if $\mu(X_{1},X_{2})=0,$ and therefore:
$(p_{00},p_{01},p_{10},p_{11})=(\theta_{1}\theta_{2},\theta_{1}(1-\theta_{2}),(1-\theta_{1})\theta_{2},(1-\theta_{1})(1-\theta_{2}))\,,$ (10)
as expected. For two dependent Bernoulli random variables, that is when $\theta_{12}\neq\theta_{1}\theta_{2},$ or equivalently when $\mu(X_{1},X_{2})\neq 0,$ without loss of generality assume that $\theta_{1}\leq\theta_{2}.$ The Fréchet-Hoeffding upper bound (3) is attained with dependence parameter $\theta_{12}=\min\{\theta_{1},\theta_{2}\}=\theta_{1}$ and therefore $\mu(X_{1},X_{2})=+1$ with:
$(p_{00},p_{01},p_{10},p_{11})=(\theta_{1},0,\theta_{2}-\theta_{1},1-\theta_{2})\,.$ (11)
Only the particular case $\theta_{1}=\theta_{2}=\theta$ translates into:
$(p_{00},p_{01},p_{10},p_{11})=(\theta,0,0,1-\theta)\,,$ (12)
that is $\mathbb{P}(X_{1}=X_{2})=p_{00}+p_{11}=1.$ Equivalently, if $\theta_{1}\neq\theta_{2}$ then it is definitely not possible to get $\mathbb{P}(X_{1}=X_{2})=1.$
The Fréchet-Hoeffding lower bound (3) is reached with dependence parameter $\theta_{12}=\max\{\theta_{1}+\theta_{2}-1,0\}$ which applied to formulas (8) translates into:
$(p_{00},p_{01},p_{10},p_{11})=\begin{cases}\,(\theta_{1}+\theta_{2}-1,1-\theta_{2},1-\theta_{1},0)&\mbox{if }\theta_{1}+\theta_{2}>1\,,\\ \,(0,\theta_{1},\theta_{2},1-\theta_{1}-\theta_{2})&\mbox{if }\theta_{1}+\theta_{2}\leq 1\,,\end{cases}$ (13)
with $\mu(X_{1},X_{2})=-1,$ from where it is clear that $\mathbb{P}(X_{1}=1-X_{2})=1$ if and only if $\theta_{1}+\theta_{2}=1.$
In summary, formulas (8) allow to define the family $\mathcal{B}_{2}(\theta_{1},\theta_{2},\theta_{12})$ of all bivariate distributions with Bernoulli univariate marginals, which is in fact also known as a Fréchet class of the type $\mathcal{F}(F_{1},F_{2}),$ see for example [9]. The bivariate joint probability mass function $p_{ij}=\mathbb{P}(X_{1}=i,X_{2}=j)$ for $i$ and $j$ in $\{0,1\}$ can be specified in two ways:
  • a)
    choosing freely the two marginal parameters $10<\theta_{1}<$ 1 and $10<\theta_{2}<$ 1 and then choosing $\theta_{12}$ such that (5) holds, or
  • b)
    choosing freely the dependence parameter $10<\theta_{12}<$ 1 and then choosing the marginal parameters $10<\theta_{1}<$ 1 and $10<\theta_{2}<$ 1 such that (5) holds. Equivalently, instead of choosing the dependence parameter value, we may choose a given value of the dependence measure defined by [3] and then obtain the value for $\theta_{12}$ from (9).
Behind the above cases a) and b) is what is known as a compatibility problem, in this particular case for the Fréchet class $\mathcal{F}(F_{1},F_{2}):$ eventhough the three parameters are the probabilities $\theta_{1}=\mathbb{P}(X_{1}=0),$ $\theta_{2}=\mathbb{P}(X_{2}=0),$ and $\theta_{12}=\mathbb{P}(X_{1}=0,X_{2}=0),$ not all values in the $[0,1]$ interval are simultaneously admissible (or compatible) for such parameters, since (5) must hold, see Figure 1.
Figure 1: Bivariate Bernoulli. Left: Compatible values for the bivariate dependence parameter $\theta_{12}$ (red interval), given a specific pair of values $(\theta_{1}^{*},\theta_{2}^{*})$ for the univariate marginal parameters. Right: Compatible values for the marginal univariate parameters $(\theta_{1},\theta_{2})$ (red triangle region), given a specific value $\theta_{12}^{*}$ for the bivariate dependence parameter.
4 Trivariate Bernoulli
Let $(X_{1},X_{2},X_{3})$ be a vector of Bernoulli random variables with (univariate marginal) parameters $1-\theta_{r}$ where $10<\theta_{r}<$ 1 as explained in the previous section, but now $r\in\{1,2,3\}.$ Accordingly to (1) the domain of the underlying trivariate subcopula is the set $\mbox{Dom}\,S^{(3)}=\{0,\theta_{1},1\}\times\{0,\theta_{2},1\}\times\{0,\theta_{3},1\}$ and:
$S^{(3)}(u,v,w)=\begin{cases}\,1&\mbox{ if }(u,v,w)=(1,1,1),\\ \,\theta_{1}&\mbox{ if }(u,v,w)=(\theta_{1},1,1),\\ \,\theta_{2}&\mbox{ if }(u,v,w)=(1,\theta_{2},1),\\ \,\theta_{3}&\mbox{ if }(u,v,w)=(1,1,\theta_{3}),\\ \,\theta_{12}&\mbox{ if }(u,v,w)=(\theta_{1},\theta_{2},1),\\ \,\theta_{13}&\mbox{ if }(u,v,w)=(\theta_{1},1,\theta_{3}),\\ \,\theta_{23}&\mbox{ if }(u,v,w)=(1,\theta_{2},\theta_{3}),\\ \,\theta_{123}&\mbox{ if }(u,v,w)=(\theta_{1},\theta_{2},\theta_{3}),\\ \,0&\mbox{ elsewhere.}\end{cases}$ (14)
Recall from Section 2 that for a trivariate subcopula $S^{(3)}$ there are $\binom{3}{2}=3$ bivariate marginal subcopulas, which in this case have bivariate dependence parameters $\theta_{12},$ $\theta_{13},$ and $\theta_{23},$ respectively, plus a trivariate dependence parameter $\theta_{123}.$ The bivariate dependence parameters must satisfy (5), that is:
$\max\{\theta_{r}+\theta_{t}-1,0\}\,\leq\,\theta_{rt}\,\leq\,\min\{\theta_{r},\theta_{t}\}\,,\quad 1\leq r<t\leq 3\,.$ (15)
The trivariate dependence parameter $\theta_{123}$ must satisfy the Fréchet-Hoeffding bounds (3), which in this case translates into:
$\max\{\theta_{1}+\theta_{2}+\theta_{3}-2,0\}\,\leq\,\theta_{123}\,\leq\,\min\{\theta_{1},\theta_{2},\theta_{3}\}\,,$ (16)
but it will be discussed later that (16) is a necessary but could not be a sufficient condition for an admisible value for $\theta_{123},$ due to a compatibility problem, similarly as mentioned in the bivariate case. Since $\mbox{Ran}\,(X_{1},X_{2},X_{3})=\{(i,j,k):i,j,k\in\{0,1\}\}$ let:
$p_{ijk}:=\mathbb{P}(X_{1}=i,X_{2}=j,X_{3}=k)\,,\qquad i,j,k\in\{0,1\},$ (17)
where necessarily $0\leq p_{ijk}\leq 1$ and $\sum_{i,j,k}p_{i,j,k}=1,$ and therefore only seven of the $p_{ijk}$ need to be specified, since the eighth is just $1$ minus the sum of the other seven. Applying (1) and (14) to calculate each $p_{ijk}$ in terms of the joint distribution function $F_{123}(X_{1}\leq i,X_{2}\leq j,X_{3}\leq k):$
$p_{ijk}$ $=$ $F_{123}(i,j,k)-F_{123}(i,j,k-1)-F_{123}(i,j-1,k)\ldots$
$-F_{123}(i-1,j,k)+F_{123}(i,j-1,k-1)+F_{123}(i-1,j,k-1)\ldots$
$+F_{123}(i-1,j-1,k)-F_{123}(i-1,j-1,k-1)\,,$
$=$ $S^{(3)}\left(F_{1}(i),F_{2}(j),F_{3}(k)\right)-S^{(3)}\left(F_{1}(i),F_{2}(j),F_{3}(k-1)\right)\ldots$
$-S^{(3)}\left(F_{1}(i),F_{2}(j-1),F_{3}(k)\right)-S^{(3)}\left(F_{1}(i-1),F_{2}(j),F_{3}(k)\right)\ldots$
$+S^{(3)}\left(F_{1}(i),F_{2}(j-1),F_{3}(k-1)\right)+S^{(3)}\left(F_{1}(i-1),F_{2}(j),F_{3}(k-1)\right)\ldots$
$+S^{(3)}\left(F_{1}(i-1),F_{2}(j-1),F_{3}(k)\right)-S^{(3)}\left(F_{1}(i-1),F_{2}(j-1),F_{3}(k-1)\right)\,,$
and therefore:
$p_{000}$ $=$ $\theta_{123}\,,$ (19)
$p_{001}$ $=$ $\theta_{12}-\theta_{123}\,,$ (20)
$p_{010}$ $=$ $\theta_{13}-\theta_{123}\,,$ (21)
$p_{100}$ $=$ $\theta_{23}-\theta_{123}\,,$ (22)
$p_{011}$ $=$ $\theta_{1}-\theta_{12}-\theta_{13}+\theta_{123}\,,$ (23)
$p_{101}$ $=$ $\theta_{2}-\theta_{12}-\theta_{23}+\theta_{123}\,,$ (24)
$p_{110}$ $=$ $\theta_{3}-\theta_{13}-\theta_{23}+\theta_{123}\,,$ (25)
$p_{111}$ $=$ $1-\theta_{1}-\theta_{2}-\theta_{3}+\theta_{12}+\theta_{13}+\theta_{23}-\theta_{123}\,,$ (26)
where (26) is equal to $1$ minus the sum of (19) through (25). From (20), (21), (22), and (16), it is clear that:
$\theta_{123}\,\leq\,\min\{\theta_{12},\theta_{13},\theta_{23}\}\,\leq\,\min\{\theta_{1},\theta_{2},\theta_{3}\}$ (27)
a sharper upper bound for $\theta_{123}$ than the Fréchet-Hoeffding bounds (16), as a consequence of the compatibility problem for the Fréchet class $\mathcal{F}\{F_{12},F_{13},F_{23}\},$ but all the inequalities that arise from (19) to (26) have to be analyzed under the restrictions $0\leq p_{ijk}\leq 1$ to solve the full compatibility problem of all the parameters. It is straightforward to obtain expressions for the $\theta$ parameters in terms of the $p_{ijk}$ probabilities:
$\theta_{123}$ $=$ $p_{000}=\mathbb{P}(X_{1}=0,X_{2}=0,X_{3}=0)\,,$ (28)
$\theta_{12}$ $=$ $p_{000}+p_{001}=\mathbb{P}(X_{1}=0,X_{2}=0)\,,$ (29)
$\theta_{13}$ $=$ $p_{000}+p_{010}=\mathbb{P}(X_{1}=0,X_{3}=0)\,,$ (30)
$\theta_{23}$ $=$ $p_{000}+p_{100}=\mathbb{P}(X_{2}=0,X_{3}=0)\,,$ (31)
$\theta_{1}$ $=$ $p_{000}+p_{001}+p_{010}+p_{011}=\mathbb{P}(X_{1}=0)\,,$ (32)
$\theta_{2}$ $=$ $p_{000}+p_{001}+p_{100}+p_{101}=\mathbb{P}(X_{2}=0)\,,$ (33)
$\theta_{3}$ $=$ $p_{000}+p_{010}+p_{100}+p_{110}=\mathbb{P}(X_{3}=0)\,.$ (34)
We may follow the order in (28) through (34) to specify all the parameters such that the full compatibility of the Fréchet class $\mathcal{F}\{F_{12},F_{13},F_{23},F_{1},F_{2},F_{3}\}$ is guaranteed:
Algorithm 1
  • Step 0:
    Choose $0\leq p_{000}\leq 1$ and set the trivariate dependence parameter $\theta_{123}=p_{000}\,.$
  • Step 1:
    Choose $0\leq p_{001}\leq 1,$ $0\leq p_{010}\leq 1,$ and $0\leq p_{100}\leq 1,$ such that $p_{000}+p_{001}+p_{010}+p_{100}\leq 1,$ and calculate the bivariate dependence parameters $\theta_{rt}$ accordingly to (29), (30), and (31).
  • Step 2:
    Choose $0\leq p_{011}\leq 1,$ $0\leq p_{101}\leq 1,$ and $0\leq p_{110}\leq 1,$ such that $p_{000}+p_{001}+p_{010}+p_{100}+p_{011}+p_{101}+p_{110}\leq 1,$ and calculate the univariate (marginal) parameters $\theta_{r}$ accordingly to (32), (33), and (34).
  • Step 3:
    Calculate $p_{111}=1-p_{000}-p_{001}-p_{010}-p_{011}-p_{100}-p_{101}-p_{110}\,.$
Finally, we may apply (9) to calculate the bivariate dependence measures $\mu_{rt}=\mu(X_{r},X_{t})$ for $1\leq r<t\leq 3,$ and following the same idea:
$\mu_{123}=\begin{cases}\frac{\theta_{123}-\theta_{1}\theta_{2}\theta_{3}}{\min\{\theta_{1},\theta_{2},\theta_{3}\}-\theta_{1}\theta_{2}\theta_{3}}&\mbox{ if}\theta_{123}\geq\theta_{1}\theta_{2}\theta_{3}\,,\\ {}&{}\textstyle\\ \frac{\theta_{123}-\theta_{1}\theta_{2}\theta_{3}}{\theta_{1}\theta_{2}\theta_{3}-\max\{\theta_{1}+\theta_{2}+\theta_{3}-2,0\}}&\mbox{ if }\theta_{123}<\theta_{1}\theta_{2}\theta_{3}\,.\end{cases}$ (35)
where from (16) $-1\leq\mu_{123}\leq+1,$ and clearly $\mu_{123}=0$ if and only if the three random variables are independent.
Example 1. Consider $X_{1},X_{2},X_{3}$ identically distributed Bernoulli random variables with parameter $1-\theta,$ where $0<\theta<1,$ that are pairwise independent but not necessarily jointly independent. We will obtain all the compatible values for the trivariate dependence parameter $\theta_{123},$ and the resulting trivariate probability mass function (pmf). Combining the inequalities $0\leq p_{ijk}\leq 1$ (19) through (26) with $\theta_{r}=\theta$ and $\theta_{rt}=\theta^{2}$ (pairwise independence) we obtain:
$\max\{0,2\theta^{2}-\theta\}\,\leq\,\theta_{123}\,\leq\,\min\{\theta^{2},1+\theta-2\theta^{2},1-3\theta+3\theta^{2}\}$ (36)
which is equivalent to:
$\theta_{123}\in\begin{cases}\,\left[0\,,\,\theta^{2}\right]&\text{ if }\theta\leq\frac{1}{2}\\ \,\left[2\theta^{2}-\theta\,,\,1-3\theta+3\theta^{2}\right]&\text{ if }\frac{1}{2}<\theta\leq\frac{4}{5}\\ \,\left[2\theta^{2}-\theta\,,\,1+\theta-2\theta^{2}\right]&\text{ if }\frac{4}{5}<\theta\leq\frac{1+\sqrt{5}}{4}\end{cases}$ (37)
In Figure 2 it is depicted the set of all compatibility values $(\theta,\theta_{123}),$ and in Table 1 the resulting pmf in the particular case when $\theta=1/2$ and the special cases of extreme dependence ($\theta_{123}\in\{0,1/4\}$) and independence ($\theta_{123}=1/8$).
Figure 2: Compatible values $(\theta,\theta_{123})$ (red region)
in Example 1 for a Trivariate Bernoulli, with identical univariate marginals with parameter $1-\theta,$ pairwise independent, with trivariate dependence parameter $\theta_{123}.$
$\mu_{123}=-1$ $\mu_{123}=0$ $\mu_{123}=+1/3$
$(i,j,k)$ $p_{ijk}$ $\theta_{123}=0$ $\theta_{123}=1/8$ $\theta_{123}=1/4$
$(0,0,0)$ $\theta_{123}$ 00 $1/8$ $1/4$
$(0,0,1)$ $\theta^{2}-\theta_{123}$ $1/4$ $1/8$ 00
$(0,1,0)$ $\theta^{2}-\theta_{123}$ $1/4$ $1/8$ 00
$(1,0,0)$ $\theta^{2}-\theta_{123}$ $1/4$ $1/8$ 00
$(0,1,1)$ $\theta-2\theta^{2}+\theta_{123}$ 00 $1/8$ $1/4$
$(1,0,1)$ $\theta-2\theta^{2}+\theta_{123}$ 00 $1/8$ $1/4$
$(1,1,0)$ $\theta-2\theta^{2}+\theta_{123}$ 00 $1/8$ $1/4$
$(1,1,1)$ $1-3\theta+3\theta^{2}-\theta_{123}$ $1/4$ $1/8$ 00
Table 1: Probability mass function of a trivariate Bernoulli in Example 1, with identical univariate marginals with parameter $1-\theta=1/2,$ pairwise independent, and with trivariate parameter $\theta_{123}\in\{0,1/8,1/4\}$ corresponding to the lowest compatible value, independence, and the highest compatible value, respectively.
Example 2. Algorithm 1 will be applied to obtain non independent $X_{1},X_{2},X_{3}$ but identically distributed Bernoulli random variables with parameter $1-\theta,$ where $0<\theta<1,$ with pairwise conditional independence, but not pairwise independent. For the sake of simplicity we will assume that the bivariate dependence parameter is the same for all pairs, say $\beta:=\theta_{12}=\theta_{13}=\theta_{23}$ and let $\tau:=\theta_{123}$ be the trivariate dependence parameter.
  • Step 1)
    Choose $p_{000}\in[0,1]$ and set the trivariate dependence parameter $\tau=p_{000}.$
  • Step 2)
    Choose $p_{001},p_{010},p_{100}\in[0,1]$ such that:
    $p_{000}+p_{001}+p_{010}+p_{100}\leq 1$ (38)
    but from (20), (21), and (22), we get $p_{001}=\beta-\tau=p_{010}=p_{100}$ and therefore $\beta\geq\tau.$ From (38) $p_{001}+p_{010}+p_{100}\leq 1-\tau$ implies $\beta\leq(1+2\tau)/3,$ so combining altogether:
    $\tau\,\leq\,\beta\,\leq\,\frac{1+2\tau}{3}$ (39)
  • Step 3)
    Choose $p_{011},p_{101},p_{110}\in[0,1]$ such that:
    $p_{000}+p_{001}+p_{010}+p_{100}+p_{011}+p_{101}+p_{110}\leq 1$ (40)
    but from (23), (24), and (25), we get $p_{011}=\theta-2\beta+\tau=p_{101}=p_{110}$ and therefore from (40) we obtain:
    $\theta\,\leq\,\frac{1-\tau}{3}+\beta$ (41)
  • Step 4)
    Calculate $p_{111}=1-\tau-3(\beta-\tau)-3(\theta-2\beta+\tau)=1-\tau+3(\beta-\theta).$
In summary, once specified the trivariate dependence parameter $0\leq\tau\leq 1$ then we have to choose the bivariate dependence parameter $\beta$ according to (39) and the univariate parameter $\theta$ according to (41), for compatibility. If in addition we want pairwise conditional independence, say for example $X_{1}$ and $X_{2}$ conditionally independent given the value of $X_{3}:$
$X_{1}\perp X_{2}\,|\,X_{3}=k\,\,\text{ where }k\in\{0,1\}\quad\Leftrightarrow\\ \begin{cases}\mathbb{P}(X_{1}=0,X_{2}=0\,|\,X_{3}=0)=\mathbb{P}(X_{1}=0\,|\,X_{3}=0)\mathbb{P}(X_{2}=0\,|\,X_{3}=0)\\ \mathbb{P}(X_{1}=0,X_{2}=0\,|\,X_{3}=1)=\mathbb{P}(X_{1}=0\,|\,X_{3}=1)\mathbb{P}(X_{2}=0\,|\,X_{3}=1)\end{cases}$
$X_{1}\perp X_{2}\,|\,X_{3}=k$$\Leftrightarrow$ $\begin{cases}\frac{p_{000}}{\theta}=\frac{p_{000}+p_{010}}{\theta}\cdot\frac{p_{000}+p_{100}}{\theta}\\ \frac{p_{001}}{1-\theta}=\frac{p_{001}+p_{011}}{1-\theta}\cdot\frac{p_{001}+p_{101}}{1-\theta}\end{cases}$ (42)
$\Leftrightarrow$ $\begin{cases}\theta\tau=\beta^{2}\\ (1-\theta)(\beta-\theta)=(\theta-\beta)^{2}\end{cases}$
$\Leftrightarrow$ $\tau^{2}(\beta-\tau)=\beta^{3}(\beta-\tau)$
where (42) is always satisfied whenever $\beta=\tau$ which in turn would imply that $\theta=\beta$ and since $\beta\neq\theta^{2}$ nor $\tau=\theta^{3}$ for $10<\theta<$ 1 then the condition $\theta=\beta=\tau$ implies pairwise dependence and trivariate dependence but with pairwise conditional independence. In fact, in terms of the probabilities $p_{ijk}$ this implies $p_{000}=\theta,$ $p_{111}=1-\theta$ and zero probability for the remaining six cases.
5 Multivariate Bernoulli
Let $(X_{1},\ldots,X_{n})$ be a vector of $n\geq 2$ Bernoulli random variables with (univariate marginal) parameters $1-\theta_{r}$ where $0<\theta_{r}<1,$ $r\in\{1,\ldots,n\}.$ Accordingly to (1) the domain of the underlying multivariate subcopula is the set $\operatorname{Dom}\,S^{(n)}=\{0,\theta_{1},1\}\,\times\,\cdots\,\times\,\{0,\theta_{n},1\}$ and:
$S^{(n)}(u_{1},\ldots,u_{n})=\begin{cases}\,1&\mbox{ if }u_{r}=1\mbox{ for all }r\in\{1,\ldots,n\},\\ \,\theta_{s}&\mbox{ if }u_{s}=\theta_{s}\mbox{ and }u_{r}=1\mbox{ for all }r\neq s,\\ \,\theta_{rs}&\mbox{ if }r<s,u_{r}=\theta_{r},u_{s}=\theta_{s},\mbox{ and }u_{t}=1\mbox{ for all }t\neq r,s,\\ \,\theta_{rst}&\mbox{ if }r<s<t,u_{r}=\theta_{r},u_{s}=\theta_{s},u_{t}=\theta_{t},u_{\ell}=1,\ell\neq r,s,t,\\ \quad\vdots&\qquad\vdots\\ \,\theta_{12\cdots n}&\mbox{ if }u_{1}=\theta_{1},\ldots,u_{n}=\theta_{n},\\ \,0&\mbox{ elsewhere.}\end{cases}$ (43)
with a total of $12^{n}-$ 1 parameters. For any ordered subset of $\{1,\ldots,n\},$ say $A=\{a_{1},\ldots,a_{m}\}$ where $m\leq n$ and if $r<s$ then $a_{r}<a_{s},$ the $m$-variate dependence parameter $\theta_{\!A}$ must satisfy the Fréchet-Hoeffding bounds (3), which in this case translates into:
[…]
Since $\mbox{Ran}\,(X_{1},\ldots,X_{n})=\{0,1\}^{n}$ let:
$p_{i_{1}\ldots i_{n}}:=\mathbb{P}(X_{1}=i_{1},\ldots,X_{n}=i_{n})\,,\qquad i_{r}\in\{0,1\},$ (45)
where necessarily $0\leq p_{i_{1}\ldots i_{n}}\leq 1$ and $\sum\cdots\sum p_{i_{1}\ldots i_{n}}=1,$ and therefore just $12^{n}-$ 1 of the $p_{i_{1}\ldots i_{n}}$ need to be specified, since the $2^{n}$-th is just $1$ minus the sum of the other $12^{n}-$ 1 . Then expressions for the $\theta_{\!A}$ parameters in terms of (45) probabilities are obtained by:
$\theta_{\!A}\,=\,\!\!\!\!\!\!\!\!\!\!\!\!\sum_{\qquad\qquad i_{r}\,\notin\,A}\!\!\!\!\!\!\!\!\!\!\!\!\cdots\sum p_{i_{1}\ldots i_{n}}$ (46)
which can be specified according to the following algorithm to guarantee full compatibility of the dependence parameters, with the conventional notation that if $D$ is a set of numbers then $\sum D$ is equivalent to $\sum_{d\in D}d\,:$
Algorithm 2
  • Step 0:
    Choose $0\leq p_{0\cdots 0}\leq 1,$ and define the singleton set $B_{0}:=\{p_{0\cdots 0}\}.$ Set the $n$-variate dependence parameter $\theta_{1\cdots n}=p_{0\cdots 0}.$
  • Step 1:
    Choose $B_{1}:=\{0\leq p_{i_{1}\ldots i_{n}}\leq 1:\sum i_{r}=1\}$ such that $\sum B_{0}\cup B_{1}\leq 1,$ and set the $(n-1)$-variate parameters $\theta_{\!A}$ accordingly to (46).
  • Step 2:
    Choose $B_{2}:=\{0\leq p_{i_{1}\ldots i_{n}}\leq 1:\sum i_{r}=2\}$ such that $\sum B_{0}\cup B_{1}\cup B_{2}\leq 1,$ and set the $(n-2)$-variate parameters $\theta_{\!A}$ accordingly to (46).
  • Step $k$:
[…]
  • Step $n$:
[…]
Finally, we may calculate $m$-variate dependence measures for $2\leq m\leq n$ as follows:
$\mu_{A}=\begin{cases}\frac{\theta_{\!A}-\prod_{r\in A}\theta_{r}}{\min\{\theta_{a}:a\in A\}-\prod_{r\in A}\theta_{r}}&\mbox{ if }\theta_{\!A}\geq\prod_{r\in A}\theta_{r}\,,\\ {}&{}\textstyle\\ \frac{\theta_{\!A}-\prod_{r\in A}\theta_{r}}{\prod_{r\in A}\theta_{r}-\max\{\sum_{a\in A}\theta_{a}-m+1,0\}}&\mbox{ if }\theta_{\!A}<\prod_{r\in A}\theta_{r}\,.\end{cases}$ (47)
where from (44) $-1\leq\mu_{A}\leq+1,$ and clearly $\mu_{A}=0$ if and only if the random variables indexed by $A$ are independent.
5.1 Bayesian inference
If $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{m}\}$ is a size $m$ observed random sample from a Multivariate Bernoulli random vector $\mathbf{X}=(X_{1},\ldots,X_{n})$ where $\mathbf{x}_{j}=(x_{j,1},\ldots,x_{j,n})$ for $j\in\{1,\ldots,m\},$ we can make statistical inferences about the $2^{n}$ parameters $p_{i_{1}\ldots i_{n}}.$ For a more simplified notation, following [17], since $(i_{1},\ldots,i_{n})\in\{0,1\}^{n}$ we may consider it as a $n$-positions binary number and apply the one-to-one mapping conversion to a positive integer. For $r\in\{1,\ldots,2^{n}-1\}$ let $B(r-1)=[i_{1}(r),\ldots,i_{n}(r)]$ be the $n$-positions binary representation of $r-1$, so that:
$p_{r}=\mathbb{P}[X_{1}=i_{1}(r),\ldots,X_{n}=i_{n}(r)]$ (48)
and:
[…]
Then the likelihood function associated to the unknown parameters $p_{1},\ldots,p_{2^{n}-1}$ is given by:
[…]
where:
[…]
Clearly (50) has as conjugate prior distribution a $12^{n}-$ 1 dimensional Dirichlet (Di) distribution with $2^{n}$ hyperparameters $\alpha_{r}>0,$ see for example [1], and therefore the posterior distribution is given by:
$\pi(\mathbf{p}\,|\,\mathbf{x})\sim\text{Di}_{\,2^{n}-1}(\alpha_{1}+n_{1},\ldots,\alpha_{2^{n}-1}+n_{2^{n}-1},\alpha_{2^{n}}+m-\sum_{r\,=\,1}^{2^{n}-1}n_{r})$ (52)
where $\mathbf{p}=(p_{1},\ldots,p_{2^{n}-1})$ and $\mathbf{x}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{m}).$ With (52) we can also make posterior inferences for the $\theta_{\!A}$ parameters through (46), and for the $\mu_{A}$ parameters through (47).
5.2 Computational implementation
The Julia programming language [2] has been used for the implementation of all the previous results, and the code is available in a public repository at [4] for full reproducibility. The main functions implemented:
  • MBerDep Calculates dependence parameters $\theta_{A}$ and measures $\mu_{A}$ given the parameters $p_{1},\ldots,p_{2^{n}}$ of a $n$-dimensional multivariate Bernoulli distribution.
  • MBerMargin Calculates dependence parameters and measures for the $m$-dimensional marginal distribution of a $n$-dimensional multivariate Bernoulli distribution, where $1\leq m\leq n.$
  • MBerCond Calculates conditional probabilities, dependence parameters and measures from a $n$-dimensional multivariate Bernoulli distribution, given the values for a subset of random variables.
  • MBerSim Simulates a random sample from a $n$-dimensional multivariate Bernoulli distribution.
  • MBerBayes Posterior Dirichlet model for the parameters of a $n$-dimensional multivariate Bernoulli distribution given an observed random sample.
  • MBerInf Posterior point and interval estimation for the probability parameters, and the dependence parameters and measures, given an observed sample from a multivariate Bernoulli distribution, and a prior value. For dimension $10$ or higher calculations may take more than an hour. In such case you may consider using MBerEst just for point estimations.
Example 1 revisited. If we choose $\theta=0.6$ then according to (37) we must choose $\theta_{123}\in[0.12,0.28]$ where $\theta^{3}=0.216$ belongs to such interval and would represent the case of joint independence. Let’s use $\theta_{123}=0.15$ and applying formulas (19) through (26) we may obtain the $p_{ijk}$ joint probabilities $p_{000}=0.15,$ $p_{001}=0.21,$ $p_{010}=0.21,$ $p_{011}=0.03,$ $p_{100}=0.21,$ $p_{101}=0.03,$ $p_{110}=0.03,$ and $p_{111}=0.13.$ With such values me define in Julia the vector:
julia> pp = [0.15,0.21,0.21,0.03,0.21,0.03,0.03,0.13];
julia> X = MBerDep(pp);
And then verify the we get the expected dependence parameters and measures:
julia> [X.dparam.idx X.dparam.value]
72 Matrix{Any}:
 [1, 2, 3]  0.15
 [1, 2]     0.36
 [1, 3]     0.36
 [2, 3]     0.36
 [1]        0.6
 [2]        0.6
 [3]        0.6
from where we get that the bivariate dependence parameters $\theta_{12}=0.36=\theta_{13}=\theta_{23}=\theta^{2}$ as expected (pairwise independence) but certainly not joint independence since $\theta_{123}=0.15\neq 0.216=\theta^{3}.$ For the dependence measures we get:
julia> [X.dmeas.idx X.dmeas.value]
42 Matrix{Any}:
 [1, 2, 3]  -0.305556
 [1, 2]      0.0
 [1, 3]      0.0
 [2, 3]      0.0
with pairwise dependencies $\mu_{ij}$ all equal to zero (as expected) but with a 3-variate negative dependence $\mu_{123}=-0.306$ as expected, since we chose a value $\theta_{123}=0.15$ below the value that would represent independence ($\theta^{3}=0.216$).
Example 3. A trivariate Bernoulli distribution will be defined applying Algorithm 1 from Section 4 as follows:
julia> begin
           p000 = 0.1
           p001, p010, p100 = 0.2, 0.1, 0.05
           p011, p101, p110 = 0.2, 0.15, 0.1
           p111 = 1 - sum([p000, p001, p010, p011, p100, p101, p110])
           pp = [p000, p001, p010, p011, p100, p101, p110, p111]
       end
8-element Vector{Float64}:
 0.1
 0.2
 0.1
 0.2
 0.05
 0.15
 0.1
 0.09999999999999987
julia> X = MBerDep(pp); X.nrvs
3
where X.nrvs is the corresponding dimension of the random vector. Then we simulate a random sample of size $3,000$ and make point and interval estimations about the probability parameters, and the dependence parameters and measures, and a comparison to the theoretical values. In the estimation matrices the third column is always a median posterior point estimation, and the second and fourth columns are the extreme points of a $99\%$ probability posterior interval estimation:
julia> begin
           Random.seed!(1234) # for reproducibility
           simX = MBerSim(pp, 3_000);
           infer = MBerInf(simX, prior = 1/2, nsim = 10_000, probint = 0.99);
           infer.probs # estimated probabilities
       end
84 Matrix{Any}:
 [0, 0, 0]  0.0861693  0.0996407  0.1143
 [0, 0, 1]  0.176657   0.19481    0.21376
 [0, 1, 0]  0.087351   0.100971   0.115828
 [0, 1, 1]  0.175131   0.193508   0.212438
 [1, 0, 0]  0.0332592  0.0420294  0.0520198
 [1, 0, 1]  0.139136   0.155535   0.173265
 [1, 1, 0]  0.090522   0.104274   0.119274
 [1, 1, 1]  0.0945739  0.108592   0.1238
julia> [X.binprob.idx X.binprob.value] # theoretical probabilities
82 Matrix{Any}:
 [0, 0, 0]  0.1
 [0, 0, 1]  0.2
 [0, 1, 0]  0.1
 [0, 1, 1]  0.2
 [1, 0, 0]  0.05
 [1, 0, 1]  0.15
 [1, 1, 0]  0.1
 [1, 1, 1]  0.1
julia> infer.dparam # estimated dependence parameters
74 Matrix{Any}:
 [1, 2, 3]  0.0861693  0.0996407  0.1143
 [1, 2]     0.273408   0.294522   0.316158
 [1, 3]     0.182337   0.200692   0.219919
 [2, 3]     0.125958   0.141746   0.158641
 [1]        0.566172   0.589215   0.612288
 [2]        0.468764   0.492305   0.515974
 [3]        0.325026   0.347202   0.369748
julia> [X.dparam.idx X.dparam.value] # theoretical dependence parameters
72 Matrix{Any}:
 [1, 2, 3]  0.1
 [1, 2]     0.3
 [1, 3]     0.2
 [2, 3]     0.15
 [1]        0.6
 [2]        0.5
 [3]        0.35
julia> infer.dmeas # estimated dependence measures
44 Matrix{Any}:
 [1, 2, 3]  -0.113144   -0.0104273   0.039112
 [1, 2]     -0.0338129   0.0222768   0.0793568
 [1, 3]     -0.072454   -0.0188205   0.0504544
 [2, 3]     -0.23491    -0.170452   -0.105142
julia> [X.dmeas.idx X.dmeas.value] # theoretical dependence measures
42 Matrix{Any}:
 [1, 2, 3]  -0.047619
 [1, 2]      0.0
 [1, 3]     -0.047619
 [2, 3]     -0.142857
Example 4. Data from official statistics about COVID-19 in Mexico during the first year of the pandemic, available for download from [4] as a covid2020.csv file, with 2.15 million patient registries and 15 binary variables measured (among many others available).
begin
    df = CSV.read("covid2020.csv", DataFrame)
    show(describe(df), allrows = true)
    data = zeros(Int, size(df))
    for c in 1:ncol(df)
        data[:, c] = df[:, c]
    end
    println()
    data
end
15x7 DataFrame
 Row | variable        mean        min    median   max    nmissing  eltype
     | Symbol          Float64     Int64  Float64  Int64  Int64     DataType
   1 | SEXO            0.501803        0      1.0      1         0  Int64
   2 | TIPO_PACIENTE   0.196705        0      0.0      1         0  Int64
   3 | DIABETES        0.1348          0      0.0      1         0  Int64
   4 | EPOC            0.0113235       0      0.0      1         0  Int64
   5 | INMUSUPR        0.00843405      0      0.0      1         0  Int64
   6 | HIPERTENSION    0.174666        0      0.0      1         0  Int64
   7 | CARDIOVASCULAR  0.0158313       0      0.0      1         0  Int64
   8 | OBESIDAD        0.144864        0      0.0      1         0  Int64
   9 | RENAL_CRONICA   0.0152094       0      0.0      1         0  Int64
  10 | TABAQUISMO      0.0735159       0      0.0      1         0  Int64
  11 | e00             0.0502618       0      0.0      1         0  Int64
  12 | e20             0.38682         0      0.0      1         0  Int64
  13 | e40             0.434831        0      0.0      1         0  Int64
  14 | e65             0.128087        0      0.0      1         0  Int64
  15 | MUERTE          0.101051        0      0.0      1         0  Int64
2154719x15 Matrix{Int64}:
 0  0  1  0  0  0  0  0  0  0  0  0  1  0  0
 1  0  0  0  0  0  0  0  0  0  0  1  0  0  0
 0  0  1  0  0  0  0  0  0  0  0  0  1  0  0
 1  1  1  1  0  0  1  1  0  0  0  0  0  1  1
 0  1  1  0  0  1  0  1  0  0  0  0  1  0  1
 1  1  0  0  0  0  0  1  0  0  0  0  0  1  1
With the following description of the variables:
  1. 1.
    SEXO = $1$ for male, 00 for female.
  2. 2.
    TIPO_PACIENTE = $1$ for hospitalized, 00 non-hospitalized.
  3. 3.
    DIABETES = $1$ for diabetic, 00 if not.
  4. 4.
    EPOC = $1$ for chronic obstructive pulmonary disease (COPD), 00 if not.
  5. 5.
    INMUSUPR = $1$ for immunosuppression, 00 if not.
  6. 6.
    HIPERTENSION = $1$ for hypertension, 00 if not.
  7. 7.
    CARDIOVASCULAR = $1$ for cardiovascular disease, 00 if not.
  8. 8.
    OBESIDAD = $1$ for obesity, 00 if not.
  9. 9.
    RENAL_CRONICA = $1$ for chronic kidney disease (CKD), 00 if not.
  10. 10.
    TABAQUISMO = $1$ for smoker, 00 for non-smoker.
  11. 11.
    e00 = $1$ for ages between 00 and $19$, 00 if not.
  12. 12.
    e20 = $1$ for ages between $20$ and $39$, 00 if not.
  13. 13.
    e40 = $1$ for ages between $40$ and $64$, 00 if not.
  14. 14.
    e65 = $1$ for ages between $65$ and older, 00 if not.
  15. 15.
    MUERTE = $1$ for death, 00 if survived.
In this case $n=15$ and therefore there are $7672^{15}-1=32,$ 767 probabilities to estimate, $12^{15}-$ 1 parameters to estimate, and $162^{15}-$ 16 multivariate dependencies to estimate (pairwise, three-wise, up to 15-wise).
@time estim = MBerEst(data); # 5 minutes approx
If we are interested in the 20 largest values for dependence parameters that include variable 15 (death) then:
begin
    iord = sortperm(estim.dmeas.value, rev = true)
    midx = estim.dmeas.idx[iord]
    m = estim.dmeas.value[iord]
    iMue = findall(x -> 15 in x, midx)
    [midx[iMue] m[iMue]][1:20, :]
end
202 Matrix{Any}:
 [2, 15]              0.902144
 [2, 4, 15]           0.859008
 [2, 9, 15]           0.858144
 [2, 5, 15]           0.855051
 [2, 4, 9, 15]        0.831122
 [2, 7, 15]           0.831098
 [2, 5, 9, 15]        0.828303
 [2, 4, 5, 15]        0.827159
 [2, 7, 9, 15]        0.809365
 [2, 4, 5, 9, 15]     0.808968
 [2, 4, 7, 15]        0.808805
 [2, 5, 7, 15]        0.803182
 [2, 4, 7, 9, 15]     0.793962
 [2, 5, 7, 9, 15]     0.789702
 [2, 4, 5, 7, 15]     0.787861
 [2, 4, 5, 7, 9, 15]  0.778416
 [2, 4, 9, 14, 15]    0.659362
 [2, 4, 14, 15]       0.658259
 [2, 14, 15]          0.656512
 [2, 9, 14, 15]       0.65601
Therefore from the 20 highest values for multivariate dependence that always include death we are able to identify in all cases variable 2 (hospitalization), and in several cases variables 4, 9, 5, 7 and 14 (COPD, CKD, immunosuppression, cardiovascular disease, and age 65 or older). In this data the mortality rate was $10.1\%$ but conditioning on the mentioned variables it increases up to $66.3\%:$
begin
    pd = mean(data[:, 15])
    estim2 = MBerEst(data[:, [2,4,5,7,9,14,15]])
    pdcond = MBerCond(estim2.binprob.value, [7], [1,2,3,4,5,6],
                      [1,1,1,1,1,1]).binprob.dic[[1]]
    println("P(death) = ", pd)
    println("P(death|hospital, COPD, immunosup, cardio, CKD,
             age 65+) = ", pdcond)
end
P(death) = 0.10105122756145929
P(death|hosp,COPD,immunosup,cardio,CKD,65+) = 0.6632653061224489
6 Conclusions
In this paper, a novel subcopula-based characterization of dependence for the Multivariate Bernoulli Distribution (MBD) has been introduced. By leveraging Sklar’s theorem, it has been shown that subcopulas can effectively capture the dependence structure between Bernoulli random variables, extending the utility of subcopula theory to binary data. The proposed approach provides a flexible framework for studying multivariate binary data by separating the marginal distributions from the joint dependence structure.
Explicit formulas and bounds for both bivariate and multivariate Bernoulli distributions in terms of subcopulas were derived, allowing for a more granular understanding of the dependence parameters. In particular, dependence measures were obtained for bivariate and trivariate Bernoulli distributions, and proposed a general method to extend these results to higher dimensions. This work’s findings confirm that subcopulas are a powerful tool for modeling and quantifying dependence in the MBD of all orders, not just pairwise.
The proposed methodology not only advances the theoretical understanding of subcopulas but also opens up new possibilities for practical applications in fields dealing with multivariate binary data. Potential applications include genetics, epidemiology, and social sciences, where understanding the dependence structure between binary variables is essential. Moreover, the bayesian inference framework presented offers a robust tool for parameter estimation and uncertainty quantification, enhancing the practical implementation of the proposed approach.
Computational implementations of the MBD are quite scarce, so the set of functions provided in subsection 5.2 and downloadable from [4] is a contribution that may help in analyzing and simulating multivariate binary data. Future research could explore further generalizations of the subcopula approach to other types of discrete data and investigate its performance in applied contexts. Additionally, incorporating covariates or other forms of heterogeneity into the subcopula framework may provide more nuanced models, particularly for data sets with more complex dependence structures.

### v2

@@
 Abstract
-By applying Sklar’s theorem to the Multivariate Bernoulli Distribution (MBD), it is proposed a framework that decouples the marginal distributions from the dependence structure, providing a clearer understanding of how binary variables interact.
+By applying Sklar’s theorem to the Multivariate Bernoulli Distribution (MBD), this paper proposes a framework to decouple marginal distributions from the dependence structure, clarifying interactions among binary variables.
 Explicit formulas are derived under the MBD using subcopulas to introduce dependence measures for interactions of all orders, not just pairwise.
 A Bayesian inference approach is also applied to estimate the parameters of the MBD, offering practical tools for parameter estimation and dependence analysis in real-world applications.
-The results obtained contribute to the application of subcopulas of multivariate binary data, with a real data example of comorbidities in COVID-19 patients.
-Keywords: Sklar theorem, subcopula, dependence, Bernoulli, binary data.
+The results obtained contribute to the application of subcopulas of multivariate binary data, with real data examples.
+keywords:
+Sklar’s theorem, subcopula, dependence, Bernoulli, binary data
 1 Introduction
@@
 This is especially true when working with multivariate data where traditional correlation measures fall short due to the binary nature of the variables and the dimensionality.
-A representation of the MBD that provides an alternative to the traditional log-linear model for binary variables was proposed by [17] using the concept of Kronecker product from matrix calculus, in terms of $12^{n}-$ 1 parameters, where $n$ is the dimension of a random vector with MBD, but where the dependency vector of parameters are, in fact, non standardized central moments.
+A representation of the MBD that provides an alternative to the traditional log-linear model for binary variables was proposed by [1] using the concept of Kronecker product from matrix calculus, in terms of $2^{n}-1$ parameters, where $n$ is the dimension of a random vector with MBD, but where the dependency vector of parameters is, in fact, non standardized central moments.
 In [2] several flexible methods for simulating random binary sequences with fixed marginal distributions and specified degrees of association between the variables are discussed, but only pairwise dependencies are considered.
 An algorithm is derived in [3] for generating systems of correlated binary data, but allowing for the specification of just pairwise correlations within each system.
-A simple method to characterize multivariate Bernoulli variables with given means is investigated in [5] but only pairwise are correlations are discussed.
+A simple method to characterize multivariate Bernoulli variables with given means is investigated in [4] but only pairwise correlations are discussed.
 Copulas, and more generally subcopulas, provide a straightforward approach to describing the dependence structure in multivariate settings by decoupling the marginal distributions from the joint distribution.
 Introduced by Sklar’s theorem [5], subcopulas extend copula theory to cases where the marginals are not necessarily continuous, making them particularly useful for multivariate discrete data.
-This paper focuses on leveraging subcopulas to characterize the dependence within the MBD, offering a framework to better understand and quantify dependence of all orders (pariwise, three-wise, four-wise, etc.)
-Building upon previous work in copula theory and dependence measures, the present work introduces a formal subcopula-based characterization of the dependence structure within the MBD, and illustrates its usefulness to tackle the compatibility problem of building multivariate models in terms of lower dimension marginals.
+In [6] it is discussed that many of the convenient properties of copulas do not carry over from the continuous to the discrete case, because the underlying copula is not unique, but according to [5] there is a unique subcopula, and that is the reason why this paper focuses on leveraging subcopulas to characterize the dependence within the MBD, offering a framework to better understand and quantify dependence of all orders (pairwise, three-wise, four-wise, etc.)
+The present work introduces a formal subcopula-based characterization of the dependence structure within the MBD, and illustrates its usefulness to tackle the compatibility problem of building multivariate models in terms of lower dimension marginals.
 This work provides a systematic approach for analyzing the joint dependence of multivariate binary data, with potential applications in fields ranging from genetics to econometrics where binary outcomes are prevalent.
@@
 2 Subcopulas and dependence
-Sklar’s Theorem [15] proves that for any $n-$dimensional vector of random variables $(X_{1},\ldots,X_{n})$ there exists a functional link $S^{(n)}$ between its probability joint distribution $F_{X_{1},\ldots,X_{n}}(x_{1},\ldots,x_{n})=\mathbb{P}(X_{1}\leq x_{1},\ldots,X_{n}\leq x_{n})$ and its corresponding univariate marginal distributions $F_{X_{i}}(x_{i})=\mathbb{P}(X_{i}\leq x_{i}),$ $i\in\{1,\ldots,n\}:$
+Sklar’s Theorem [5] proves that for any $n$-dimensional vector of random variables $(X_{1},\ldots,X_{n})$ there exists a functional link $S^{(n)}$ between its joint probability distribution $F_{X_{1},\ldots,X_{n}}(x_{1},\ldots,x_{n})=\mathbb{P}(X_{1}\leq x_{1},\ldots,X_{n}\leq x_{n})$ and its corresponding univariate marginal distributions $F_{X_{i}}(x_{i})=\mathbb{P}(X_{i}\leq x_{i}),$ $i\in\{1,\ldots,n\}:$
 $F_{X_{1},\ldots,X_{n}}(x_{1},\ldots,x_{n})\,=\,S^{(n)}\left(F_{X_{1}}(x_{1}),\ldots,F_{X_{n}}(x_{n})\right)\,$, (1)
-for any $(x_{1},\ldots,x_{n})\in\overline{\mathbb{R}}^{n},$ and where the function $S^{(n)}:D_{1}\times\cdots\times D_{n}\rightarrow[0,1]$ is unique, with $D_{i}=\mbox{Ran}\,F_{X_{i}}$ for $i\in\{1,\ldots,n\},$ which implies $\{0,1\}\subseteq D_{i}\subseteq[0,1].$ Such link $S^{(n)}$ is called a subcopula function, and in the particular case where every $D_{i}=[0,1]$ then it is called a copula function, which would be the case when all the random variables are continuous.
-If at least one of the random variables is not continuous then the domain of the underlying subcopula $S^{(n)}$ is a proper subset of the unit hypercube $[0,1]^{n}.$ A formal definition for subcopula functions and their properties are discussed in detail in [13] and [12].
-Recalling that the random variables $X_{1},\ldots,X_{n}$ are independent if and only if their probability joint distribution is equal to the product of its univariate marginals, then as an immediate consequence of (1) the unique subcopula function under independence is:
+for any $(x_{1},\ldots,x_{n})\in\overline{\mathbb{R}}^{n},$ and where the function $S^{(n)}:D_{1}\times\cdots\times D_{n}\rightarrow[0,1]$ is unique, with $D_{i}=\mbox{Ran}\,F_{X_{i}}=\{F_{X_{i}}(x):x\in\overline{\mathbb{R}}\}$ for $i\in\{1,\ldots,n\},$ which implies $\{0,1\}\subseteq D_{i}\subseteq[0,1].$ Such link $S^{(n)}$ is called a subcopula function, and in the particular case where every $D_{i}=[0,1]$ then it is called a copula function, which would be the case when all the random variables are continuous.
+If at least one of the random variables is not continuous then the domain of the underlying subcopula $S^{(n)}$ is a proper subset of the unit hypercube $0,1]^{n}.$ A formal definition for subcopula functions and their properties are discussed in detail in [8] and [9].
+Recalling that the random variables $X_{1},\ldots,X_{n}$ are mutually independent if and only if their probability joint distribution is equal to the product of its univariate marginals, then as an immediate consequence of (1) the unique subcopula function under independence is:
 $\Pi^{(n)}(u_{1},\ldots,u_{n}):=u_{1}\cdots u_{n}\,,\qquad(u_{1},\ldots,u_{n})\in D_{1}\times\cdots\times D_{n}\,.$ (2)
-Moreover, as a consequence of the Fréchet-Hoeffding bounds for joint distributions (see [6] and [8]) and applying (1) we may obtain bounds for any subcopula $S^{(n)}:$
+Moreover, as a consequence of the Fréchet-Hoeffding bounds for joint distributions (see [10] and [11]) and, applying (1), we may obtain bounds for any subcopula $S^{(n)}:$
 $W^{(n)}(u_{1},\ldots,u_{n})\,\leq\,S^{(n)}(u_{1},\ldots,u_{n})\,\leq\,M^{(n)}(u_{1},\ldots,u_{n})\,$, (3)
 where $W^{(n)}(u_{1},\ldots,u_{n}):=\max\{u_{1}+\cdots+u_{n}-n+1,0\}$ and $M^{(n)}(u_{1},\ldots,u_{n}):=\min\{u_{1},\ldots,u_{n}\}.$ The upper bound $M^{(n)}$ is always a subcopula, but the lower bound $W^{(n)}$ is subcopula only for $n=2,$ though it is still the best possible lower bound for $n>2,$ see [9].
-From any $n-$dimensional subcopula $S^{(n)}(u_{1},\ldots,u_{n})$ and $2\leq k<n$ it is possible to obtain $\binom{n}{k}$ $k-$dimensional marginal subcopulas $S^{(k)}(v_{1},\ldots,v_{k})$ defined in terms of $S^{(n)}(u_{1},\ldots,u_{n})$ where $k$ of its entries are equal to exactly one of each $\{v_{1},\ldots,v_{k}\}$ and the remaining $n-k$ are equal to $1$
-Since the univariate marginal distributions $F_{X_{i}}(x_{i})=\mathbb{P}(X_{i}\leq x_{i})$ have no information about how each random variable $X_{i}$ interacts with others, then as a consequence of (1) all the information about the dependence among the random variables $X_{1},\ldots,X_{n}$ is contained in their unique underlying subcopula $S^{(n)},$ and therefore any attempt to measure degrees or intensity of dependence should extract information from $S^{(n)}.$ For the particular case of bivariate copulas (when both random variables are continuous) several copula-based measures have been studied, such as the concordance measures by Spearman [16] or Kendall [10], or dependence measures by Schweizer-Wolff [14] or Hoeffding [8].
-For the case of proper subcopulas (that are not copulas) a subcopula dependence measure has been proposed by Erdely [3].
+From any $n$-dimensional subcopula $S^{(n)}(u_{1},\ldots,u_{n})$ and $2\leq k<n$ it is possible to obtain $\binom{n}{k}$ $k$-dimensional marginal subcopulas $S^{(k)}(v_{1},\ldots,v_{k})$ defined in terms of $S^{(n)}(u_{1},\ldots,u_{n})$ where $k$ of its entries are equal to exactly one of each $\{v_{1},\ldots,v_{k}\}$ and the remaining nkn-k are equal to $1.$
+Since the univariate marginal distributions $F_{X_{i}}(x_{i})=\mathbb{P}(X_{i}\leq x_{i})$ have no information about how each random variable $X_{i}$ interacts with others, then as a consequence of (1) all the information about the dependence among the random variables $X_{1},\ldots,X_{n}$ is contained in their unique underlying subcopula $S^{(n)},$ and therefore any attempt to measure degrees or intensity of dependence should extract information from $S^{(n)}.$
+For the particular case of bivariate copulas (when both random variables are continuous) several copula-based measures have been studied, such as the concordance measures by Spearman [12] or Kendall [13], or dependence measures by Schweizer-Wolff [14] or Hoeffding [11].
+But when at least one of the variables is discrete, as pointed out by [6], an identifiability issue arises, and therefore copula-based dependence or concordance measures cannot be adapted in this case since it is not possible to report a unique value.
+This motivated the work by [15] where for the case of proper subcopulas (that are not copulas) a subcopula dependence measure has been proposed.
+For the case of a vector $(X,Y)$ of arbitrary type random variables with underlying and unique subcopula $S$ it is defined:
+$d(S):=\sup_{\text{Dom}\,S}\{S-\Pi_{S}\}-\sup_{\text{Dom}\,S}\{\Pi_{S}-S\}\,.$ (4)
+where for any $(u,v)\in\text{Dom}\,S=\text{Ran}\,F_{X}\times\text{Ran}\,F_{Y}$ the following subcopulas are defined as $W_{S}(u,v):=\max\{u+v-1,0\},$ $M_{S}(u,v):=\min\{u,v\},$ and $\Pi_{S}(u,v)=uv,$ and then:
+[…]
+where $d(M_{S})$ and $d(W_{S})$ is formula (4) applied to $M_{S}$ and $W_{S},$ respectively.
+In (4) it is calculated the largest difference between the unique underlying subcopula $S$ and the subcopula $\Pi_{S}$ that represents independence, and this value is used in (5) to calculate a proportion with respect to the largest possible differences, according to the Frécht-Hoeffding bounds for subcopulas.
+In fact, (4) and (5) may be similarly applied to any subcopula in higher dimensions, using (2) and (3).
 3 Bivariate Bernoulli
+Let $(X_{1},X_{2})$ be a vector of Bernoulli random variables with (univariate marginal) parameters $1-\theta_{r}$ where $0<\theta_{r}<1,$ that is $\mathbb{P}(X_{r}=0)=\theta_{r}$ and $\mathbb{P}(X_{r}=1)=1-\theta_{r}$ for $r\in\{1,2\}.$ Then the univariate marginal distribution functions are given by $F_{r}(x)=\mathbb{P}(X_{r}\leq x)=\theta_{r}\mathbb{I}_{\{0\,\leq\,x\,<\,1\}}+\mathbb{I}_{\{x\,\geq\,1\}},$ where $\mathbb{I}_{A}$ stands for the indicator (or characteristic) function of the set $A,$ and therefore $\mbox{Ran}\,F_{r}=\{0,\theta_{r},1\}.$
+According to (1) the domain of the underlying subcopula in this case is the set $\mbox{Dom}\,S^{(2)}=\{0,\theta_{1},1\}\times\{0,\theta_{2},1\}.$ From the formal definition of bivariate subcopula (see [9], for example) it always satisfies $S^{(2)}(u,0)=0=S^{(2)}(0,v),$ $S^{(2)}(u,1)=u,$ and $S^{(2)}(1,v)=v,$ and therefore:
 […]
-Accordingly to (1) the domain of the underlying subcopula in this case is the set $\mbox{Dom}\,S^{(2)}=\{0,\theta_{1},1\}\times\{0,\theta_{2},1\}.$ From the formal definition of bivariate subcopula (see [12], for example) it always satisfies $S^{(2)}(u,0)=0=S^{(2)}(0,v),$ $S^{(2)}(u,1)=u,$ and $S^{(2)}(1,v)=v,$ and therefore:
-$S^{(2)}(u,v)=\begin{cases}\,1&\mbox{ if }(u,v)=(1,1),\\ \,\theta_{1}&\mbox{ if }(u,v)=(\theta_{1},1),\\ \,\theta_{2}&\mbox{ if }(u,v)=(1,\theta_{2}),\\ \,\theta_{12}&\mbox{ if }(u,v)=(\theta_{1},\theta_{2}),\\ \,0&\mbox{ elsewhere,}\end{cases}$ (4)
 where $\theta_{12}$ is any value satisfying:
@@
 Since $\mbox{Ran}\,(X,Y)=\{(0,0),(0,1),(1,0),(1,1)\}$ let $\{p_{00},p_{01},p_{10},p_{11}\}$ be the joint point probabilities:
-$p_{ij}:=\mathbb{P}(X_{1}=i,X_{2}=j)\,,\qquad i,j\in\{0,1\},$ (6)
+$p_{ij}:=\mathbb{P}(X_{1}=i,X_{2}=j)\,,\qquad i,j\in\{0,1\},$
 where necessarily $0\leq p_{ij}\leq 1$ and $p_{00}+p_{01}+p_{10}+p_{11}=1,$ and therefore only three of the $p_{ij}$ values need to be specified, since the fourth is just $1$ minus the sum of the other three.
 Applying (1) and (6) to calculate each $p_{ij}$ in terms of the joint distribution function $F_{12}(i,j)=\mathbb{P}(X_{1}\leq i,X_{2}\leq j):$
-$p_{ij}$ $=$ $F_{12}(i,j)-F_{12}(i,j-1)-F_{12}(i-1,j)+F_{12}(i-1,j-1)\,,$ (7)
+$p_{ij}$ $=$ $F_{12}(i,j)-F_{12}(i,j-1)-F_{12}(i-1,j)+F_{12}(i-1,j-1)\,$,
 $=$ $S^{(2)}\left(F_{1}(i),F_{2}(j)\right)-S^{(2)}\left(F_{1}(i),F_{2}(j-1)\right)-S^{(2)}\left(F_{1}(i-1),F_{2}(j)\right)\ldots$
@@
 $p_{11}$ $=$ $1-\theta_{1}-\theta_{2}+\theta_{12}\,.$
-Formulas (8) allow to fully characterize the family $\mathcal{B}_{2}(\theta_{1},\theta_{2},\theta_{12})$ of all bivariate Bernoulli distributions in terms of the marginal parameters $10<\theta_{1}<$ 1 and $0<\theta_{2}<1,$ and a bivariate dependence parameter $\theta_{12}$ that must satisfy $\max\{\theta_{1}+\theta_{2}-1,0\}\leq\theta_{12}\leq\min\{\theta_{1},\theta_{2}\}.$ The subcopula based dependence measure proposed by [3] for $\mathcal{B}_{2}(\theta_{1},\theta_{2},\theta_{12})$ becomes:
+Formulas (8) allow us to fully characterize the family of all bivariate Bernoulli distributions in terms of the marginal parameters $0<\theta_{1}<1$ and $0<\theta_{2}<1,$ and a bivariate dependence parameter $\theta_{12}$ that must satisfy $\max\{\theta_{1}+\theta_{2}-1,0\}\leq\theta_{12}\leq\min\{\theta_{1},\theta_{2}\},$ that we may denote as $\mathcal{B}_{2}(\theta_{1},\theta_{2},\theta_{12}).$ The subcopula based dependence measure proposed by [15] for $\mathcal{B}_{2}(\theta_{1},\theta_{2},\theta_{12})$ becomes:
 $\mu(X_{1},X_{2})=\begin{cases}\frac{\theta_{12}-\theta_{1}\theta_{2}}{\min\{\theta_{1},\theta_{2}\}-\theta_{1}\theta_{2}}&\mbox{ if }\theta_{12}\geq\theta_{1}\theta_{2}\,,\\ {}&{}\textstyle\\ \frac{\theta_{12}-\theta_{1}\theta_{2}}{\theta_{1}\theta_{2}-\max\{\theta_{1}+\theta_{2}-1,0\}}&\mbox{ if }\theta_{12}<\theta_{1}\theta_{2}\,.\end{cases}$ (9)
 Two Bernoulli random variables are independent if and only if $\theta_{12}=\theta_{1}\theta_{2}$ as an immediate consequence of (2), or equivalently if and only if $\mu(X_{1},X_{2})=0,$ and therefore:
-$(p_{00},p_{01},p_{10},p_{11})=(\theta_{1}\theta_{2},\theta_{1}(1-\theta_{2}),(1-\theta_{1})\theta_{2},(1-\theta_{1})(1-\theta_{2}))\,,$ (10)
+$(p_{00},p_{01},p_{10},p_{11})=(\theta_{1}\theta_{2},\theta_{1}(1-\theta_{2}),(1-\theta_{1})\theta_{2},(1-\theta_{1})(1-\theta_{2}))\,$,
 as expected.
 For two dependent Bernoulli random variables, that is when $\theta_{12}\neq\theta_{1}\theta_{2},$ or equivalently when $\mu(X_{1},X_{2})\neq 0,$ without loss of generality assume that $\theta_{1}\leq\theta_{2}.$ The Fréchet-Hoeffding upper bound (3) is attained with dependence parameter $\theta_{12}=\min\{\theta_{1},\theta_{2}\}=\theta_{1}$ and therefore $\mu(X_{1},X_{2})=+1$ with:
-$(p_{00},p_{01},p_{10},p_{11})=(\theta_{1},0,\theta_{2}-\theta_{1},1-\theta_{2})\,.$ (11)
+$(p_{00},p_{01},p_{10},p_{11})=(\theta_{1},0,\theta_{2}-\theta_{1},1-\theta_{2})\,.$
 Only the particular case $\theta_{1}=\theta_{2}=\theta$ translates into:
-$(p_{00},p_{01},p_{10},p_{11})=(\theta,0,0,1-\theta)\,,$ (12)
+$(p_{00},p_{01},p_{10},p_{11})=(\theta,0,0,1-\theta)\,$,
 that is $\mathbb{P}(X_{1}=X_{2})=p_{00}+p_{11}=1.$ Equivalently, if $\theta_{1}\neq\theta_{2}$ then it is definitely not possible to get $\mathbb{P}(X_{1}=X_{2})=1.$
 The Fréchet-Hoeffding lower bound (3) is reached with dependence parameter $\theta_{12}=\max\{\theta_{1}+\theta_{2}-1,0\}$ which applied to formulas (8) translates into:
-$(p_{00},p_{01},p_{10},p_{11})=\begin{cases}\,(\theta_{1}+\theta_{2}-1,1-\theta_{2},1-\theta_{1},0)&\mbox{if }\theta_{1}+\theta_{2}>1\,,\\ \,(0,\theta_{1},\theta_{2},1-\theta_{1}-\theta_{2})&\mbox{if }\theta_{1}+\theta_{2}\leq 1\,,\end{cases}$ (13)
+[…]
 with $\mu(X_{1},X_{2})=-1,$ from where it is clear that $\mathbb{P}(X_{1}=1-X_{2})=1$ if and only if $\theta_{1}+\theta_{2}=1.$
 In summary, formulas (8) allow to define the family $\mathcal{B}_{2}(\theta_{1},\theta_{2},\theta_{12})$ of all bivariate distributions with Bernoulli univariate marginals, which is in fact also known as a Fréchet class of the type $\mathcal{F}(F_{1},F_{2}),$ see for example [16].
+A Fréchet class is a set of all multivariate distribution functions that share the same given marginal distributions.
+Formally, let $F_{1},F_{2},\ldots,F_{d}$ be univariate cummulative distribution functions.
+The Fréchet class associated with $(F_{1},\ldots,F_{d})$ is defined as:
+$\mathcal{F}(F_{1},\ldots,F_{d}):=\{H:H\text{ is a $d$-dimensional distribution with marginals }F_{1},\ldots,F_{d}\}$
 The bivariate joint probability mass function $p_{ij}=\mathbb{P}(X_{1}=i,X_{2}=j)$ for $i$ and $j$ in $\{0,1\}$ can be specified in two ways:
   • a)
-    choosing freely the two marginal parameters $10<\theta_{1}<$ 1 and $10<\theta_{2}<$ 1 and then choosing $\theta_{12}$ such that (5) holds, or
+    choosing freely the two marginal parameters $0<\theta_{1}<1$ and $0<\theta_{2}<1$ and then choosing $\theta_{12}$ such that (7) holds, or
   • b)
-    choosing freely the dependence parameter $10<\theta_{12}<$ 1 and then choosing the marginal parameters $10<\theta_{1}<$ 1 and $10<\theta_{2}<$ 1 such that (5) holds.
+    choosing freely the dependence parameter $0<\theta_{12}<1$ and then choosing the marginal parameters $0<\theta_{1}<1$ and $0<\theta_{2}<1$ such that (7) holds.
 Equivalently, instead of choosing the dependence parameter value, we may choose a given value of the dependence measure defined by [15] and then obtain the value for $\theta_{12}$ from (9).
-Behind the above cases a) and b) is what is known as a compatibility problem, in this particular case for the Fréchet class $\mathcal{F}(F_{1},F_{2}):$ eventhough the three parameters are the probabilities $\theta_{1}=\mathbb{P}(X_{1}=0),$ $\theta_{2}=\mathbb{P}(X_{2}=0),$ and $\theta_{12}=\mathbb{P}(X_{1}=0,X_{2}=0),$ not all values in the $[0,1]$ interval are simultaneously admissible (or compatible) for such parameters, since (5) must hold, see Figure 1.
+Behind the above cases a) and b) is what is known as a compatibility problem, in this particular case for the Fréchet class $\mathcal{F}(F_{1},F_{2}):$ even though the three parameters are the probabilities $\theta_{1}=\mathbb{P}(X_{1}=0),$ $\theta_{2}=\mathbb{P}(X_{2}=0),$ and $\theta_{12}=\mathbb{P}(X_{1}=0,X_{2}=0),$ not all values in the $[0,1]$ interval are simultaneously admissible (or compatible) for such parameters, since (7) must hold, see Figure 1:
+  • a)
+    Figure 1 Left: Specific values are chosen for $\theta_{1}=\theta_{1}^{*}$ and $\theta_{2}=\theta_{2}^{*}$ and from (7) all the admissible values for $\theta_{12}\in\big{[}\max\{\theta_{1}^{*}+\theta_{2}^{*}-1,0\}\,,\,\min\{\theta_{1}^{*},\theta_{2}^{*}\}\big{]}.$ If the chosen point $(\theta_{1}^{*},\theta_{2}^{*})$ is below the solid gray identity line $\theta_{2}=\theta_{1}$ then $\min\{\theta_{1}^{*},\theta_{2}^{*}\}=\theta_{1}^{*},$ which is projected to the horizontal axis through such identity line (horizontally through black solid line, then vertically through a violet solid line to point $(0,\theta_{1}^{*})$), to get the upper bound for $\theta_{12}.$ For the lower bound, since $\theta_{1}+\theta_{2}-1\geq 0$ if and only if $\theta_{2}\geq 1-\theta_{1},$ then if the chosen point $(\theta_{1}^{*},\theta_{2}^{*})$ is above the solid gray line $\theta_{2}=1-\theta_{1},$ we may build a solid black line $\theta_{2}=-\theta_{1}+(\theta_{1}^{*}+\theta_{2}^{*})$ that includes the point $(\theta_{1}^{*},\theta_{2}^{*})$ and that is parallel to the solid gray line $\theta_{2}=1-\theta_{1},$ so that $\theta_{2}=1$ if and only if $\theta_{1}=\theta_{1}^{*}+\theta_{2}^{*}-1,$ which is the corresponding lower bound for $\theta_{12},$ and so the point $(\theta_{1}^{*}+\theta_{2}^{*}-1,1)$ is projected vertically down to $(0,\theta_{1}^{*}+\theta_{2}^{*}-1).$ So the red thick line represents the interval $\theta_{1}^{*}+\theta_{2}^{*}-1\,,\,\theta_{1}^{*}\,]$ of admissible values for $\theta_{12}$ in this graphical example, given chosen values $\theta_{1}=\theta_{1}^{*}$ and $\theta_{2}=\theta_{2}^{*}.$
+  • b)
+    Figure 1 Right: In this case what is chosen is a specific value $\theta_{12}=\theta_{12}^{*}$ and now will characterize the set of all points $(\theta_{1},\theta_{2})$ that are compatible with such fixed value $\theta_{12}^{*}$ according to (7), that is to solve for $\theta_{1}$ and $\theta_{2}$ these two inequalities:
+    $\max\{\theta_{1}+\theta_{2}-1,0\}$ $\leq$ $\theta_{12}^{*}\,$,
+    $\min\{\theta_{1},\theta_{2}\}$ $\geq$ $\theta_{12}^{*}\,$,
+    which in turn are equivalent to the following inequalities:
+    $\theta_{2}$ $\leq$ $\min\{1,-\theta_{1}+\theta_{12}^{*}+1\}\,$,
+[…]
+    where these last two inequalities represent the red triangle region of admissible values for $(\theta_{1},\theta_{2})$ given a fixed value $\theta_{12}=\theta_{12}^{*}.$
 Figure 1: Bivariate Bernoulli.
@@
 4 Trivariate Bernoulli
-Let $(X_{1},X_{2},X_{3})$ be a vector of Bernoulli random variables with (univariate marginal) parameters $1-\theta_{r}$ where $10<\theta_{r}<$ 1 as explained in the previous section, but now $r\in\{1,2,3\}.$ Accordingly to (1) the domain of the underlying trivariate subcopula is the set $\mbox{Dom}\,S^{(3)}=\{0,\theta_{1},1\}\times\{0,\theta_{2},1\}\times\{0,\theta_{3},1\}$ and:
-$S^{(3)}(u,v,w)=\begin{cases}\,1&\mbox{ if }(u,v,w)=(1,1,1),\\ \,\theta_{1}&\mbox{ if }(u,v,w)=(\theta_{1},1,1),\\ \,\theta_{2}&\mbox{ if }(u,v,w)=(1,\theta_{2},1),\\ \,\theta_{3}&\mbox{ if }(u,v,w)=(1,1,\theta_{3}),\\ \,\theta_{12}&\mbox{ if }(u,v,w)=(\theta_{1},\theta_{2},1),\\ \,\theta_{13}&\mbox{ if }(u,v,w)=(\theta_{1},1,\theta_{3}),\\ \,\theta_{23}&\mbox{ if }(u,v,w)=(1,\theta_{2},\theta_{3}),\\ \,\theta_{123}&\mbox{ if }(u,v,w)=(\theta_{1},\theta_{2},\theta_{3}),\\ \,0&\mbox{ elsewhere.}\end{cases}$ (14)
+Let $(X_{1},X_{2},X_{3})$ be a vector of Bernoulli random variables with (univariate marginal) parameters $1-\theta_{r}$ where $0<\theta_{r}<1$ as explained in the previous section, but now $r\in\{1,2,3\}.$ According to (1) the domain of the underlying trivariate subcopula is the set $\mbox{Dom}\,S^{(3)}=\{0,\theta_{1},1\}\times\{0,\theta_{2},1\}\times\{0,\theta_{3},1\}$ and:
+[…]
 Recall from Section 2 that for a trivariate subcopula $S^{(3)}$ there are $\binom{3}{2}=3$ bivariate marginal subcopulas, which in this case have bivariate dependence parameters $\theta_{12},$ $\theta_{13},$ and $\theta_{23},$ respectively, plus a trivariate dependence parameter $\theta_{123}.$ The bivariate dependence parameters must satisfy (7), that is:
-$\max\{\theta_{r}+\theta_{t}-1,0\}\,\leq\,\theta_{rt}\,\leq\,\min\{\theta_{r},\theta_{t}\}\,,\quad 1\leq r<t\leq 3\,.$ (15)
+$\max\{\theta_{r}+\theta_{t}-1,0\}\,\leq\,\theta_{rt}\,\leq\,\min\{\theta_{r},\theta_{t}\}\,,\quad 1\leq r<t\leq 3\,.$
 The trivariate dependence parameter $\theta_{123}$ must satisfy the Fréchet-Hoeffding bounds (3), which in this case translates into:
 $\max\{\theta_{1}+\theta_{2}+\theta_{3}-2,0\}\,\leq\,\theta_{123}\,\leq\,\min\{\theta_{1},\theta_{2},\theta_{3}\}\,$, (11)
-but it will be discussed later that (16) is a necessary but could not be a sufficient condition for an admisible value for $\theta_{123},$ due to a compatibility problem, similarly as mentioned in the bivariate case.
+but it will be discussed later that (11) is a necessary but could not be a sufficient condition for an admissible value for $\theta_{123},$ due to a compatibility problem, similarly as mentioned in the bivariate case.
 Since $\mbox{Ran}\,(X_{1},X_{2},X_{3})=\{(i,j,k):i,j,k\in\{0,1\}\}$ let:
-$p_{ijk}:=\mathbb{P}(X_{1}=i,X_{2}=j,X_{3}=k)\,,\qquad i,j,k\in\{0,1\},$ (17)
+$p_{ijk}:=\mathbb{P}(X_{1}=i,X_{2}=j,X_{3}=k)\,,\qquad i,j,k\in\{0,1\},$
 where necessarily $0\leq p_{ijk}\leq 1$ and $\sum_{i,j,k}p_{i,j,k}=1,$ and therefore only seven of the $p_{ijk}$ need to be specified, since the eighth is just $1$ minus the sum of the other seven.
@@
 where (19) is equal to $1$ minus the sum of (12) through (18). From (13), (14), (15), and (11), it is clear that:
-$\theta_{123}\,\leq\,\min\{\theta_{12},\theta_{13},\theta_{23}\}\,\leq\,\min\{\theta_{1},\theta_{2},\theta_{3}\}$ (27)
+$\theta_{123}\,\leq\,\min\{\theta_{12},\theta_{13},\theta_{23}\}\,\leq\,\min\{\theta_{1},\theta_{2},\theta_{3}\}$
 a sharper upper bound for $\theta_{123}$ than the Fréchet-Hoeffding bounds (11), as a consequence of the compatibility problem for the Fréchet class $\mathcal{F}\{F_{12},F_{13},F_{23}\},$ but all the inequalities that arise from (12) to (19) have to be analyzed under the restrictions $0\leq p_{ijk}\leq 1$ to solve the full compatibility problem of all the parameters.
@@
 We may follow the order in (20) through (26) to specify all the parameters such that the full compatibility of the Fréchet class $\mathcal{F}\{F_{12},F_{13},F_{23},F_{1},F_{2},F_{3}\}$ is guaranteed:
-Algorithm 1
-  • Step 0:
-    Choose $0\leq p_{000}\leq 1$ and set the trivariate dependence parameter $\theta_{123}=p_{000}\,.$
-  • Step 1:
-    Choose $0\leq p_{001}\leq 1,$ $0\leq p_{010}\leq 1,$ and $0\leq p_{100}\leq 1,$ such that $p_{000}+p_{001}+p_{010}+p_{100}\leq 1,$ and calculate the bivariate dependence parameters $\theta_{rt}$ accordingly to (29), (30), and (31).
-  • Step 2:
-    Choose $0\leq p_{011}\leq 1,$ $0\leq p_{101}\leq 1,$ and $0\leq p_{110}\leq 1,$ such that $p_{000}+p_{001}+p_{010}+p_{100}+p_{011}+p_{101}+p_{110}\leq 1,$ and calculate the univariate (marginal) parameters $\theta_{r}$ accordingly to (32), (33), and (34).
-  • Step 3:
-    Calculate $p_{111}=1-p_{000}-p_{001}-p_{010}-p_{011}-p_{100}-p_{101}-p_{110}\,.$
+Parameter Selection Procedure 1
+  • Step 1: Choose $0\leq p_{000}\leq 1$ and set the trivariate dependence parameter $\theta_{123}=p_{000}\,.$
+  • Step 2: Choose $0\leq p_{001}\leq 1,$ $0\leq p_{010}\leq 1,$ and $0\leq p_{100}\leq 1,$ such that $p_{000}+p_{001}+p_{010}+p_{100}\leq 1,$ and calculate the bivariate dependence parameters $\theta_{rt}$ according to (21), (22), and (23).
+  • Step 3: Choose $0\leq p_{011}\leq 1,$ $0\leq p_{101}\leq 1,$ and $0\leq p_{110}\leq 1,$ such that $p_{000}+p_{001}+p_{010}+p_{100}+p_{011}+p_{101}+p_{110}\leq 1,$ and calculate the univariate (marginal) parameters $\theta_{r}$ according to (24), (25), and (26).
+  • Step 4: Calculate $p_{111}=1-p_{000}-p_{001}-p_{010}-p_{011}-p_{100}-p_{101}-p_{110}\,.$
 Finally, we may apply (9) to calculate the bivariate dependence measures $\mu_{rt}=\mu(X_{r},X_{t})$ for $1\leq r<t\leq 3,$ and following the same idea:
 $\mu_{123}=\begin{cases}\frac{\theta_{123}-\theta_{1}\theta_{2}\theta_{3}}{\min\{\theta_{1},\theta_{2},\theta_{3}\}-\theta_{1}\theta_{2}\theta_{3}}&\mbox{ if }\theta_{123}\geq\theta_{1}\theta_{2}\theta_{3}\,,\\ {}&{}\textstyle\\ \frac{\theta_{123}-\theta_{1}\theta_{2}\theta_{3}}{\theta_{1}\theta_{2}\theta_{3}-\max\{\theta_{1}+\theta_{2}+\theta_{3}-2,0\}}&\mbox{ if }\theta_{123}<\theta_{1}\theta_{2}\theta_{3}\,.\end{cases}$ (27)
-where from (16) $-1\leq\mu_{123}\leq+1,$ and clearly $\mu_{123}=0$ if and only if the three random variables are independent.
+where from (11) $1\leq\mu_{123}\leq+1,$ and clearly $\mu_{123}=0$ if and only if the three random variables are independent.
 Example 1.
@@
 Combining the inequalities $0\leq p_{ijk}\leq 1$ (12) through (19) with $\theta_{r}=\theta$ and $\theta_{rt}=\theta^{2}$ (pairwise independence) we obtain:
-$\max\{0,2\theta^{2}-\theta\}\,\leq\,\theta_{123}\,\leq\,\min\{\theta^{2},1+\theta-2\theta^{2},1-3\theta+3\theta^{2}\}$ (36)
+$\max\{0,2\theta^{2}-\theta\}\,\leq\,\theta_{123}\,\leq\,\min\{\theta^{2},1+\theta-2\theta^{2},1-3\theta+3\theta^{2}\}$
 which is equivalent to:
-$\theta_{123}\in\begin{cases}\,\left[0\,,\,\theta^{2}\right]&\text{ if }\theta\leq\frac{1}{2}\\ \,\left[2\theta^{2}-\theta\,,\,1-3\theta+3\theta^{2}\right]&\text{ if }\frac{1}{2}<\theta\leq\frac{4}{5}\\ \,\left[2\theta^{2}-\theta\,,\,1+\theta-2\theta^{2}\right]&\text{ if }\frac{4}{5}<\theta\leq\frac{1+\sqrt{5}}{4}\end{cases}$ (37)
+[…]
 In Figure 2 it is depicted the set of all compatibility values $(\theta,\theta_{123}),$ and in Table 1 the resulting pmf in the particular case when $\theta=1/2$ and the special cases of extreme dependence ($\theta_{123}\in\{0,1/4\}$) and independence ($\theta_{123}=1/8$).
-Figure 2: Compatible values $(\theta,\theta_{123})$ (red region)
-in Example 1 for a Trivariate Bernoulli, with identical univariate marginals with parameter $1-\theta,$ pairwise independent, with trivariate dependence parameter $\theta_{123}.$
+Figure 2: Compatible values $(\theta,\theta_{123})$ (red region) in Example 1 for a Trivariate Bernoulli, with identical univariate marginals with parameter $1-\theta,$ pairwise independent, with trivariate dependence parameter $\theta_{123}.$
 $\mu_{123}=-1$ $\mu_{123}=0$ $\mu_{123}=+1/3$
 $(i,j,k)$ $p_{ijk}$ $\theta_{123}=0$ $\theta_{123}=1/8$ $\theta_{123}=1/4$
-$(0,0,0)$ $\theta_{123}$ 00 $1/8$ $1/4$
-$(0,0,1)$ $\theta^{2}-\theta_{123}$ $1/4$ $1/8$ 00
-$(0,1,0)$ $\theta^{2}-\theta_{123}$ $1/4$ $1/8$ 00
-$(1,0,0)$ $\theta^{2}-\theta_{123}$ $1/4$ $1/8$ 00
-$(0,1,1)$ $\theta-2\theta^{2}+\theta_{123}$ 00 $1/8$ $1/4$
-$(1,0,1)$ $\theta-2\theta^{2}+\theta_{123}$ 00 $1/8$ $1/4$
-$(1,1,0)$ $\theta-2\theta^{2}+\theta_{123}$ 00 $1/8$ $1/4$
-$(1,1,1)$ $1-3\theta+3\theta^{2}-\theta_{123}$ $1/4$ $1/8$ 00
+$(0,0,0)$ $\theta_{123}$ 0 $1/8$ $1/4$
+$(0,0,1)$ $\theta^{2}-\theta_{123}$ $1/4$ $1/8$ 0
+$(0,1,0)$ $\theta^{2}-\theta_{123}$ $1/4$ $1/8$ 0
+$(1,0,0)$ $\theta^{2}-\theta_{123}$ $1/4$ $1/8$ 0
+$(0,1,1)$ $\theta-2\theta^{2}+\theta_{123}$ 0 $1/8$ $1/4$
+$(1,0,1)$ $\theta-2\theta^{2}+\theta_{123}$ 0 $1/8$ $1/4$
+$(1,1,0)$ $\theta-2\theta^{2}+\theta_{123}$ 0 $1/8$ $1/4$
+$(1,1,1)$ $1-3\theta+3\theta^{2}-\theta_{123}$ $1/4$ $1/8$ 0
 Table 1: Probability mass function of a trivariate Bernoulli in Example 1, with identical univariate marginals with parameter $1-\theta=1/2,$ pairwise independent, and with trivariate parameter $\theta_{123}\in\{0,1/8,1/4\}$ corresponding to the lowest compatible value, independence, and the highest compatible value, respectively.
@@
 For the sake of simplicity we will assume that the bivariate dependence parameter is the same for all pairs, say $\beta:=\theta_{12}=\theta_{13}=\theta_{23}$ and let $\tau:=\theta_{123}$ be the trivariate dependence parameter.
-  • Step 1)
-    Choose $p_{000}\in[0,1]$ and set the trivariate dependence parameter $\tau=p_{000}.$
-  • Step 2)
-    Choose $p_{001},p_{010},p_{100}\in[0,1]$ such that:
-    $p_{000}+p_{001}+p_{010}+p_{100}\leq 1$ (38)
+  • Step 1: Choose $p_{000}\in[0,1]$ and set the trivariate dependence parameter $\tau=p_{000}.$
+  • Step 2: Choose $p_{001},p_{010},p_{100}\in[0,1]$ such that:
+    $p_{000}+p_{001}+p_{010}+p_{100}\leq 1$
     but from (13), (14), and (15), we get $p_{001}=\beta-\tau=p_{010}=p_{100}$ and therefore $\beta\geq\tau.$ From (4) $p_{001}+p_{010}+p_{100}\leq 1-\tau$ implies $\beta\leq(1+2\tau)/3,$ so combining altogether:
     $\tau\,\leq\,\beta\,\leq\,\frac{1+2\tau}{3}$ (29)
-  • Step 3)
-    Choose $p_{011},p_{101},p_{110}\in[0,1]$ such that:
+  • Step 3: Choose $p_{011},p_{101},p_{110}\in[0,1]$ such that:
     $p_{000}+p_{001}+p_{010}+p_{100}+p_{011}+p_{101}+p_{110}\leq 1$ (30)
@@
     $\theta\,\leq\,\frac{1-\tau}{3}+\beta$ (31)
-  • Step 4)
-    Calculate $p_{111}=1-\tau-3(\beta-\tau)-3(\theta-2\beta+\tau)=1-\tau+3(\beta-\theta).$
+  • Step 4: Calculate $p_{111}=1-\tau-3(\beta-\tau)-3(\theta-2\beta+\tau)=1-\tau+3(\beta-\theta).$
 In summary, once specified the trivariate dependence parameter $0\leq\tau\leq 1$ then we have to choose the bivariate dependence parameter $\beta$ according to (29) and the univariate parameter $\theta$ according to (31), for compatibility.
@@
 $\Leftrightarrow$ $\tau^{2}(\beta-\tau)=\beta^{3}(\beta-\tau)$
-where (42) is always satisfied whenever $\beta=\tau$ which in turn would imply that $\theta=\beta$ and since $\beta\neq\theta^{2}$ nor $\tau=\theta^{3}$ for $10<\theta<$ 1 then the condition $\theta=\beta=\tau$ implies pairwise dependence and trivariate dependence but with pairwise conditional independence.
+where (32) is always satisfied whenever $\beta=\tau$ which in turn would imply that $\theta=\beta$ and since $\beta\neq\theta^{2}$ nor $\tau=\theta^{3}$ for $0<\theta<1$ then the condition $\theta=\beta=\tau$ implies pairwise dependence and trivariate dependence but with pairwise conditional independence.
 In fact, in terms of the probabilities $p_{ijk}$ this implies $p_{000}=\theta,$ $p_{111}=1-\theta$ and zero probability for the remaining six cases.
 5 Multivariate Bernoulli
-Let $(X_{1},\ldots,X_{n})$ be a vector of $n\geq 2$ Bernoulli random variables with (univariate marginal) parameters $1-\theta_{r}$ where $0<\theta_{r}<1,$ $r\in\{1,\ldots,n\}.$ Accordingly to (1) the domain of the underlying multivariate subcopula is the set $\operatorname{Dom}\,S^{(n)}=\{0,\theta_{1},1\}\,\times\,\cdots\,\times\,\{0,\theta_{n},1\}$ and:
-$S^{(n)}(u_{1},\ldots,u_{n})=\begin{cases}\,1&\mbox{ if }u_{r}=1\mbox{ for all }r\in\{1,\ldots,n\},\\ \,\theta_{s}&\mbox{ if }u_{s}=\theta_{s}\mbox{ and }u_{r}=1\mbox{ for all }r\neq s,\\ \,\theta_{rs}&\mbox{ if }r<s,u_{r}=\theta_{r},u_{s}=\theta_{s},\mbox{ and }u_{t}=1\mbox{ for all }t\neq r,s,\\ \,\theta_{rst}&\mbox{ if }r<s<t,u_{r}=\theta_{r},u_{s}=\theta_{s},u_{t}=\theta_{t},u_{\ell}=1,\ell\neq r,s,t,\\ \quad\vdots&\qquad\vdots\\ \,\theta_{12\cdots n}&\mbox{ if }u_{1}=\theta_{1},\ldots,u_{n}=\theta_{n},\\ \,0&\mbox{ elsewhere.}\end{cases}$ (43)
-with a total of $12^{n}-$ 1 parameters.
+Let $(X_{1},\ldots,X_{n})$ be a vector of $n\geq 2$ Bernoulli random variables with (univariate marginal) parameters $1-\theta_{r}$ where $0<\theta_{r}<1,$ $r\in\{1,\ldots,n\}.$ According to (1) the domain of the underlying multivariate subcopula is the set $\operatorname{Dom}\,S^{(n)}=\{0,\theta_{1},1\}\,\times\,\cdots\,\times\,\{0,\theta_{n},1\}$ and:
+[…]
+with a total of $2^{n}-1$ parameters.
 For any ordered subset of $\{1,\ldots,n\},$ say $A=\{a_{1},\ldots,a_{m}\}$ where $m\leq n$ and if $r<s$ then $a_{r}<a_{s},$ the $m$-variate dependence parameter $\theta_{\!A}$ must satisfy the Fréchet-Hoeffding bounds (3), which in this case translates into:
-[…]
+$\max\left\{\sum_{a\,\in\,A}\theta_{a}-m+1\,,\,0\right\}\,\leq\,\theta_{A}\,\leq\,\min\{\theta_{a}:a\in A\}\,.$ (33)
 Since $\mbox{Ran}\,(X_{1},\ldots,X_{n})=\{0,1\}^{n}$ let:
 $p_{i_{1}\ldots i_{n}}:=\mathbb{P}(X_{1}=i_{1},\ldots,X_{n}=i_{n})\,,\qquad i_{r}\in\{0,1\},$ (34)
-where necessarily $0\leq p_{i_{1}\ldots i_{n}}\leq 1$ and $\sum\cdots\sum p_{i_{1}\ldots i_{n}}=1,$ and therefore just $12^{n}-$ 1 of the $p_{i_{1}\ldots i_{n}}$ need to be specified, since the $2^{n}$-th is just $1$ minus the sum of the other $12^{n}-$ 1 .
-Then expressions for the $\theta_{\!A}$ parameters in terms of (45) probabilities are obtained by:
+where necessarily $0\leq p_{i_{1}\ldots i_{n}}\leq 1$ and $\sum\cdots\sum p_{i_{1}\ldots i_{n}}=1,$ and therefore just $2^{n}-1$ of the $p_{i_{1}\ldots i_{n}}$ need to be specified, since the $2^{n}$-th is just $1$ minus the sum of the other $2^{n}-1.$ Then expressions for the $\theta_{\!A}$ parameters in terms of (34) probabilities are obtained by:
 $\theta_{\!A}\,=\,\!\!\!\!\!\!\!\!\!\!\!\!\sum_{\qquad\qquad i_{r}\,\notin\,A}\!\!\!\!\!\!\!\!\!\!\!\!\cdots\sum p_{i_{1}\ldots i_{n}}$ (35)
 which can be specified according to the following algorithm to guarantee full compatibility of the dependence parameters, with the conventional notation that if $D$ is a set of numbers then $\sum D$ is equivalent to $\sum_{d\in D}d\,:$
-Algorithm 2
-  • Step 0:
-    Choose $0\leq p_{0\cdots 0}\leq 1,$ and define the singleton set $B_{0}:=\{p_{0\cdots 0}\}.$ Set the $n$-variate dependence parameter $\theta_{1\cdots n}=p_{0\cdots 0}.$
-  • Step 1:
-    Choose $B_{1}:=\{0\leq p_{i_{1}\ldots i_{n}}\leq 1:\sum i_{r}=1\}$ such that $\sum B_{0}\cup B_{1}\leq 1,$ and set the $(n-1)$-variate parameters $\theta_{\!A}$ accordingly to (46).
-  • Step 2:
-    Choose $B_{2}:=\{0\leq p_{i_{1}\ldots i_{n}}\leq 1:\sum i_{r}=2\}$ such that $\sum B_{0}\cup B_{1}\cup B_{2}\leq 1,$ and set the $(n-2)$-variate parameters $\theta_{\!A}$ accordingly to (46).
-  • Step $k$:
-[…]
-  • Step $n$:
-[…]
+Parameter Selection Procedure 2
+  • Step 1: Choose $0\leq p_{0\cdots 0}\leq 1,$ and define the singleton set $B_{0}:=\{p_{0\cdots 0}\}.$ Set the $n$-variate dependence parameter $\theta_{1\cdots n}=p_{0\cdots 0}.$
+  • Step 2: Choose $B_{1}:=\{0\leq p_{i_{1}\ldots i_{n}}\leq 1:\sum i_{r}=1\}$ such that $\sum B_{0}\cup B_{1}\leq 1,$ and set the (n1)(n-1)-variate parameters $\theta_{\!A}$ according to (35).
+  • Step 3: Choose $B_{2}:=\{0\leq p_{i_{1}\ldots i_{n}}\leq 1:\sum i_{r}=2\}$ such that $\sum B_{0}\cup B_{1}\cup B_{2}\leq 1,$ and set the (n2)(n-2)-variate parameters $\theta_{\!A}$ according to (35).
+  • Step $k+1$: Choose $B_{k}:=\{0\leq p_{i_{1}\ldots i_{n}}\leq 1:\sum i_{r}=k\}$ such that $\sum\bigcup_{t\,=\,0}^{k}B_{t}\leq 1,$ and set the (nk)(n-k)-variate parameters $\theta_{\!A}$ according to (35).
+  • Step $n+1$: Calculate $p_{1\cdots 1}=1-\sum\bigcup_{t\,=\,0}^{n-1}B_{t}\,.$
 Finally, we may calculate $m$-variate dependence measures for $2\leq m\leq n$ as follows:
 $\mu_{A}=\begin{cases}\frac{\theta_{\!A}-\prod_{r\in A}\theta_{r}}{\min\{\theta_{a}:a\in A\}-\prod_{r\in A}\theta_{r}}&\mbox{ if }\theta_{\!A}\geq\prod_{r\in A}\theta_{r}\,,\\ {}&{}\textstyle\\ \frac{\theta_{\!A}-\prod_{r\in A}\theta_{r}}{\prod_{r\in A}\theta_{r}-\max\{\sum_{a\in A}\theta_{a}-m+1,0\}}&\mbox{ if }\theta_{\!A}<\prod_{r\in A}\theta_{r}\,.\end{cases}$ (36)
-where from (44) $-1\leq\mu_{A}\leq+1,$ and clearly $\mu_{A}=0$ if and only if the random variables indexed by $A$ are independent.
+where from (33) $1\leq\mu_{A}\leq+1,$ and clearly $\mu_{A}=0$ if and only if the random variables indexed by $A$ are independent.
 5.1 Bayesian inference
+Under the Bayesian paradigm of statistics, given an observed random sample $\mathbf{y}=(y_{1},\ldots,y_{m})$ from a probability density model $f_{Y}(y\,|\,\theta),$ where $\theta\in\Theta$ is a vector of unknown parameters, any statistical inference about $\theta$ is performed through a posterior distribution for $\theta$ given the data $\mathbf{y}$ applying Bayes’s rule, that is:
+$\pi(\theta\,|\,\mathbf{y})=\frac{f(\mathbf{y}\,|\,\theta)\pi(\theta)}{\int_{\,\,\,\Theta}\!\!\!\!\!\!\cdots\!\int f(\mathbf{y}\,|\,\tilde{\theta})\pi(\tilde{\theta})\,d\tilde{\theta}}$
+where $\pi(\theta)$ is a prior distribution, and where:
+[…]
+is also known as the likelihood function.
+Whenever possible, it is quite convenient to choose as a prior distribution a conjugate family, such that the posterior belongs to that same family, and therefore it is straightforward to identify, since:
+$\pi(\theta\,|\,\mathbf{y})\,\propto\,L(\theta\,|\,\mathbf{y})\pi(\theta)\,.$
 If $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{m}\}$ is a size $m$ observed random sample from a Multivariate Bernoulli random vector $\mathbf{X}=(X_{1},\ldots,X_{n})$ where $\mathbf{x}_{j}=(x_{j,1},\ldots,x_{j,n})$ for $j\in\{1,\ldots,m\},$ we can make statistical inferences about the $2^{n}$ parameters $p_{i_{1}\ldots i_{n}}.$ For a more simplified notation, following [1], since $(i_{1},\ldots,i_{n})\in\{0,1\}^{n}$ we may consider it as a $n$-positions binary number and apply the one-to-one mapping conversion to a positive integer.
-For $r\in\{1,\ldots,2^{n}-1\}$ let $B(r-1)=[i_{1}(r),\ldots,i_{n}(r)]$ be the $n$-positions binary representation of $r-1$, so that:
-$p_{r}=\mathbb{P}[X_{1}=i_{1}(r),\ldots,X_{n}=i_{n}(r)]$ (48)
+For $r\in\{1,\ldots,2^{n}-1\}$ let $B(r-1)=[i_{1}(r),\ldots,i_{n}(r)]$ be the $n$-positions binary representation of r1r-1, so that:
+$p_{r}=\mathbb{P}[X_{1}=i_{1}(r),\ldots,X_{n}=i_{n}(r)]$
 and:
-[…]
-Then the likelihood function associated to the unknown parameters $p_{1},\ldots,p_{2^{n}-1}$ is given by:
-[…]
+$p_{2^{n}}=\mathbb{P}[X_{1}=1,X_{2}=1,\ldots,X_{n}=1]=1-\sum_{r\,=\,1}^{2^{n}-1}p_{r}\,.$
+Then the likelihood function associated with the unknown parameters $p_{1},\ldots,p_{2^{n}-1}$ is given by:
+$\mathbb{L}(p_{1},\ldots,p_{2^{n}-1}\,|\,\mathbf{x}_{1},\ldots,\mathbf{x}_{m})=\left(\prod_{r\,=\,1}^{2^{n}-1}p_{r}^{\,n_{r}}\right)\left(1-\sum_{r\,=\,1}^{2^{n}-1}p_{r}\right)^{m-\sum_{r\,=\,1}^{2^{n}-1}n_{r}}$ (37)
 where:
-[…]
-Clearly (50) has as conjugate prior distribution a $12^{n}-$ 1 dimensional Dirichlet (Di) distribution with $2^{n}$ hyperparameters $\alpha_{r}>0,$ see for example [1], and therefore the posterior distribution is given by:
+$n_{r}=\sum_{j\,=\,1}^{m}\mathbf{1}_{\{\mathbf{x}_{j}\,=\,B(r-1)\}}\,.$
+Clearly (37) has as conjugate prior distribution a $2^{n}-1$ dimensional Dirichlet (Di) distribution with $2^{n}$ hyperparameters $\alpha_{r}>0,$ see for example [17], and therefore the posterior distribution is given by:
 $\pi(\mathbf{p}\,|\,\mathbf{x})\sim\text{Di}_{\,2^{n}-1}(\alpha_{1}+n_{1},\ldots,\alpha_{2^{n}-1}+n_{2^{n}-1},\alpha_{2^{n}}+m-\sum_{r\,=\,1}^{2^{n}-1}n_{r})$ (38)
 where $\mathbf{p}=(p_{1},\ldots,p_{2^{n}-1})$ and $\mathbf{x}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{m}).$ With (38) we can also make posterior inferences for the $\theta_{\!A}$ parameters through (35), and for the $\mu_{A}$ parameters through (36).
-5.2 Computational implementation
-The Julia programming language [2] has been used for the implementation of all the previous results, and the code is available in a public repository at [4] for full reproducibility.
+6 Computational implementation
+The Julia programming language [7] has been used for the implementation of all the previous results, and the code is available in a public repository [18] to ensure full reproducibility.
+Julia offers significant advantages for high-performance numerical computing, especially for high-dimensional matrix operations and intensive statistical computations.
+Its speed and ability to handle large datasets efficiently make it particularly suitable for the implementation of the proposed methods.
+Additionally, Julia’s syntax is close to mathematical notation, facilitating the translation of theoretical concepts into computational code, see for example [19].
 The main functions implemented:
@@
 In such case you may consider using MBerEst just for point estimations.
-Example 1 revisited.
-If we choose $\theta=0.6$ then according to (37) we must choose $\theta_{123}\in[0.12,0.28]$ where $\theta^{3}=0.216$ belongs to such interval and would represent the case of joint independence.
-Let’s use $\theta_{123}=0.15$ and applying formulas (19) through (26) we may obtain the $p_{ijk}$ joint probabilities $p_{000}=0.15,$ $p_{001}=0.21,$ $p_{010}=0.21,$ $p_{011}=0.03,$ $p_{100}=0.21,$ $p_{101}=0.03,$ $p_{110}=0.03,$ and $p_{111}=0.13.$ With such values me define in Julia the vector:
-julia> pp = [0.15,0.21,0.21,0.03,0.21,0.03,0.03,0.13];
-julia> X = MBerDep(pp);
-And then verify the we get the expected dependence parameters and measures:
-julia> [X.dparam.idx X.dparam.value]
-72 Matrix{Any}:
- [1, 2, 3]  0.15
- [1, 2]     0.36
- [1, 3]     0.36
- [2, 3]     0.36
- [1]        0.6
- [2]        0.6
- [3]        0.6
-from where we get that the bivariate dependence parameters $\theta_{12}=0.36=\theta_{13}=\theta_{23}=\theta^{2}$ as expected (pairwise independence) but certainly not joint independence since $\theta_{123}=0.15\neq 0.216=\theta^{3}.$ For the dependence measures we get:
-julia> [X.dmeas.idx X.dmeas.value]
-42 Matrix{Any}:
- [1, 2, 3]  -0.305556
- [1, 2]      0.0
- [1, 3]      0.0
- [2, 3]      0.0
-with pairwise dependencies $\mu_{ij}$ all equal to zero (as expected) but with a 3-variate negative dependence $\mu_{123}=-0.306$ as expected, since we chose a value $\theta_{123}=0.15$ below the value that would represent independence ($\theta^{3}=0.216$).
-Example 3. A trivariate Bernoulli distribution will be defined applying Algorithm 1 from Section 4 as follows:
-julia> begin
+6.1 A theoretical example
+In Example 1 from Section 4, if we choose $\theta=0.6$ then according to (28) we must choose $\theta_{123}\in[0.12,0.28]$ where $\theta^{3}=0.216$ belongs to such interval and would represent the case of joint independence.
+Let’s use $\theta_{123}=0.15$ and applying formulas (12) through (19) we may obtain the $p_{ijk}$ joint probabilities $p_{000}=0.15,$ $p_{001}=0.21,$ $p_{010}=0.21,$ $p_{011}=0.03,$ $p_{100}=0.21,$ $p_{101}=0.03,$ $p_{110}=0.03,$ and $p_{111}=0.13.$ The Julia code to generate the results from Table 2 is in Appendix A.1, which basically makes use of the MBerDep function described at the beginning of Section 6.
+variables dependence parameter dependence measure
+$(X_{1},X_{2},X_{3})$ 0.15 -0.3056
+$(X_{1},X_{2})$ 0.36 0.0
+$(X_{1},X_{3})$ 0.36 0.0
+$(X_{2},X_{3})$ 0.36 0.0
+Table 2: Dependence parameters and measures corresponding to the provided values in subsection 6.1. The Julia code to generate the results is in Appendix A.1.
+All pairwise dependencies $\mu_{ij}$ in Table 2 are equal to zero (as expected) but with a 3-variate negative dependence $\mu_{123}=-0.306$ as expected, since we chose a value $\theta_{123}=0.15$ below the value that would represent independence ($\theta^{3}=0.216$).
+6.2 A simulation example
+A trivariate Bernoulli distribution will be defined applying Parameter Selection Procedure 1 from Section 4 as follows:
+  • Step 0: Choose $p_{000}=0{.}1$
+  • Step 1: Choose $(p_{001},p_{010},p_{100})=(0{.}2,0{.}1,0{.}05)$
+  • Step 2: Choose $(p_{011},p_{101},p_{110})=(0{.}2,0{.}15,0{.}1)$
+  • Step 3: Calculate $p_{111}=1-\sum_{ijk\neq 111}p_{ijk}$
+The Julia code to define a MBD, simulate data from it, and make statistical inferences is in Appendix A.2.
+A random sample of size $3,000$ is simulated, according to the above provided probabilities, using the MBerSim function described at the beginning of Section 6, which in turn uses a standard procedure to simulate from a discrete distribution with given mass probabilities.
+With the simulated data, point and interval estimations are made using the MBerInf function, for the probability parameters (Table 3), the parameters $\theta_{A}$ (Table 4), and the dependence measures $\mu_{A}$ (Table 5).
+To evaluate the finite-sample performance of the proposed estimation procedure, a Monte Carlo study was conducted and examined the empirical coverage rates of the bayesian probability intervals for the parameters of interest.
+For each scenario, data were generated under the given parameter values, and 10,000 replications were performed.
+A 99% probability interval was constructed in each replication, and the proportion of intervals containing the true parameter value was recorded.
+Ideally, this proportion should be close to the nominal level of 99%, and this is the case, since conjugate analysis was applied (Dirichlet distribution in this case), and there are no convergence issues for the same reason, since there was no need of applying MCMC algorithms for the posterior inferences.
+true value Lower Median Upper Coverage
+$(i,j,k)$ $p_{ijk}$ quantile 0.0050{.}005 quantile 0.5000{.}500 quantile 0.9950{.}995 rate (%)
+$(0,0,0)$ 0.100{.}10 0.0862 0.0996 0.1143 98.95
+$(0,0,1)$ 0.200{.}20 0.1767 0.1948 0.2138 98.98
+$(0,1,0)$ 0.100{.}10 0.0874 0.1010 0.1158 99.05
+$(1,0,0)$ 0.050{.}05 0.0333 0.0420 0.0520 99.08
+$(0,1,1)$ 0.200{.}20 0.1751 0.1935 0.2124 99.02
+$(1,0,1)$ 0.150{.}15 0.1391 0.1555 0.1733 98.92
+$(1,1,0)$ 0.100{.}10 0.0905 0.1043 0.1193 99.09
+$(1,1,1)$ 0.100{.}10 0.0946 0.1086 0.1238 99.15
+Table 3: Comparing theoretical probabilities (true values) with point estimates (median) and $99\%$ probability intervals (lower–upper) from simulated observations of the Multivariate Bernoulli Distribution defined in section 6.2.
+For the coverage rate 10,000 replications were performed in each case.
+The Julia code to generate the results is in Appendix A.2.
+true value Lower Median Upper Coverage
+$A\subseteq\{1,2,3\}$ $\theta_{\!A}$ quantile 0.0050{.}005 quantile 0.5000{.}500 quantile 0.9950{.}995 rate (%)
+$\{1,2,3\}$ 0.100{.}10 0.0862 0.0996 0.1143 98.95
+$\{1,2\}$ 0.300{.}30 0.2734 0.2945 0.3162 99.02
+$\{1,3\}$ 0.200{.}20 0.1823 0.2007 0.2199 98.97
+$\{2,3\}$ 0.150{.}15 0.1260 0.1417 0.1586 99.05
+$\{1\}$ 0.600{.}60 0.5662 0.5892 0.6123 98.93
+$\{2\}$ 0.500{.}50 0.4688 0.4923 0.5160 99.18
+$\{3\}$ 0.350{.}35 0.3250 0.3472 0.3697 98.91
+Table 4: Comparing theoretical parameters $\theta_{A}$ (true values) with point estimates (median) and $99\%$ probability intervals (lower–upper) from simulated observations of the Multivariate Bernoulli Distribution defined in section 6.2.
+For the coverage rate 10,000 replications were performed in each case.
+The Julia code to generate the results is in Appendix A.2.
+true value Lower Median Upper Coverage
+$A\subseteq\{1,2,3\}$ $\mu_{A}$ quantile 0.0050{.}005 quantile 0.5000{.}500 quantile 0.9950{.}995 rate(%)
+$\{1,2,3\}$ -0.0476 -0.1131 -0.0104 0.0391 99.15
+$\{1,2\}$ 0.0000 -0.0338 0.0223 0.0794 99.04
+$\{1,3\}$ -0.0476 -0.0725 -0.0188 0.0505 99.06
+$\{2,3\}$ -0.1429 -0.2349 -0.1705 -0.1051 99.03
+Table 5: Comparing theoretical dependence measures $\mu_{A}$ (true values) with point estimates (median) and $99\%$ probability intervals (lower–upper) from simulated observations of the Multivariate Bernoulli Distribution defined in section 6.2.
+For the coverage rate 10,000 replications were performed in each case.
+The Julia code to generate the results is in Appendix A.2.
+6.3 A real data example: Bank churn data
+Data related to churn detection in the banking sector has been analyzed by [20]: 10,000 registries where 4 out of 13 variables under consideration are binary.
+The original data may be downloaded from [21], but for the present example the 4 mentioned binary variables are available for download from [18] as a churnbindata.csv file, with the following description of the variables:
+  1. 1.
+    gender: 1=male, 0=female.
+  2. 2.
+    crcard: customer has credit card (1=yes, 0=no).
+  3. 3.
+    active: customer is an active member (1=yes, 0=no).
+  4. 4.
+    exited: customer has churned (1=yes, 0=no).
+The Julia code to fit a MBD is in Appendix A.3.
+In this case $n=4$ and therefore there are $2^{n}=16$ probabilities to estimate, $2^{n}-1=15$ parameters $\theta_{\!A}$ to estimate, and $2^{n}-(n+1)=11$ multivariate dependencies $\mu_{A}$ to estimate (pairwise, three-wise, and four-wise), where the latter are summarized in Table 6, with point and 95% probability interval estimations.
+$\mu_{A}$ lower $\mu_{A}$ median $\mu_{A}$ upper
+$A\subseteq\{1,2,3\}$ quantile 0.0250{.}025 quantile 0.5000{.}500 quantile 0.9750{.}975
+$\{1,2,3,4\}$ -0.1926 -0.1254 -0.0557
+$\{1,2,3\}$ -0.0528 0.0022 0.0198
+$\{1,2,4\}$ -0.0956 -0.0545 -0.0131
+$\{1,3,4\}$ -0.1537 -0.1252 -0.0966
+$\{2,3,4\}$ -0.1233 -0.0847 -0.0457
+$\{1,2\}$ -0.0230 0.0085 0.0361
+$\{1,3\}$ 0.0032 0.0240 0.0448
+$\{2,3\}$ -0.0502 -0.0189 0.0116
+$\{1,4\}$ -0.2267 -0.1914 -0.1559
+$\{2,4\}$ -0.0347 -0.0093 0.0373
+$\{3,4\}$ -0.3356 -0.2991 -0.2623
+Table 6: Dependencies $\mu_{A}$ of all orders for the binary variables from bank churn data in [21]. The Julia code to generate the results is in Appendix A.3.
+The main variable of interest is exited (variable 4) to be predicted by the others.
+In terms of pairwise dependencies, we may notice that crcard (variable 2) has no relevant dependence with exited since $\mu_{24}=-0{.}0093\in[-0{.}0347,0{.}0373]$ is close to zero and with probability 95% belongs to an interval containing zero and approximately centered around zero.
+We may also notice that the three-wise dependence $\mu_{134}=-0{.}1252$ is quite similar to the four-wise dependence $\mu_{1234}=-0{.1254},$ so we may discard crcard as a relevant variable to predict exited, and therefore fit a reduced trivariate MBD with the remaining variables, renumbering them as follows:
+  1. 1.
+    gender: 1=male, 0=female.
+  2. 2.
+    active: customer is an active member (1=yes, 0=no).
+  3. 3.
+    exited: customer has churned (1=yes, 0=no).
+In this case $n=3$ and therefore $2^{n}=8$ probabilities to estimate, $2^{n}-1=7$ parameters $\theta_{\!A}$ to estimate, and $2^{n}-(n+1)=4$ multivariate dependencies $\mu_{A}$ to estimate (pairwise and three-wise).
+Table 7 summarizes point and 95% probability interval estimations for the probabilities $p_{ijk},$ and tables 8 and 9 similar estimations for the $\theta_{\!A}$ parameters and $\mu_{A}$ dependencies, respectively.
+$p_{ijk}$ lower $p_{ijk}$ median $p_{ijk}$ upper
+$(i,j,k)$ quantile 0.0250{.}025 quantile 0.5000{.}500 quantile 0.9750{.}975
+$(0,0,0)$ 0.1465 0.1535 0.1606
+$(0,0,1)$ 0.0675 0.0725 0.0777
+$(0,1,0)$ 0.1793 0.1869 0.1946
+$(0,1,1)$ 0.0376 0.0414 0.0454
+$(1,0,0)$ 0.1935 0.2013 0.2092
+$(1,0,1)$ 0.0533 0.0578 0.0625
+$(1,1,0)$ 0.2459 0.2544 0.2630
+$(1,1,1)$ 0.0288 0.0321 0.0357
+Table 7: Point and 95% probability interval estimations for the probabilities $p_{ijk}$ in the trivariate MBD model for the binary variables (gender,active,exited) from bank churn data in [21].
+The Julia code to generate the results is in Appendix A.3.
+$\theta_{\!A}$ lower $\theta_{\!A}$ median $\theta_{\!A}$ upper
+$A\subseteq\{1,2,3\}$ quantile 0.0250{.}025 quantile 0.5000{.}500 quantile 0.9750{.}975
+$\{1,2,3\}$ 0.1465 0.1535 0.1606
+$\{1,2\}$ 0.2179 0.2260 0.2343
+$\{1,3\}$ 0.3311 0.3404 0.3497
+$\{2,3\}$ 0.3454 0.3548 0.3642
+$\{1\}$ 0.4445 0.4543 0.4641
+$\{2\}$ 0.4753 0.4851 0.4949
+$\{3\}$ 0.7881 0.7961 0.8039
+Table 8: Point and 95% probability interval estimations for the parameters $\theta_{\!A}$ in the trivariate MBD model for the binary variables (gender,active,exited) from bank churn data in [21].
+The Julia code to generate the results is in Appendix A.3.
+$\mu_{A}$ lower $\mu_{A}$ median $\mu_{A}$ upper
+$A\subseteq\{1,2,3\}$ quantile 0.0250{.}025 quantile 0.5000{.}500 quantile 0.9750{.}975
+$\{1,2,3\}$ -0.1537 -0.1252 -0.0966
+$\{1,2\}$ 0.0032 0.0240 0.0449
+$\{1,3\}$ -0.2270 -0.1916 -0.1561
+$\{2,3\}$ -0.3359 -0.2994 -0.2625
+Table 9: Point and 95% probability interval estimations for the dependencies $\mu_{A}$ in the trivariate MBD model for the binary variables (gender,active,exited) from bank churn data in [21].
+The Julia code to generate the results is in Appendix A.3.
+With the trivariate fitted MBD model we may proceed to calculate unconditional and conditional probabilities for the main variable of interest exited, as summarized in Table 10.
+With such probabilities we define 5 rules for prediction, and then we estimate their accuracy through simulations:
+  • Rule 1: Non conditional simulation of 10,000 observations from a Bernoulli random variable with parameter $\mathbb{P}(\text{{exited}}=1)=0{.}2039$ (case 1 in Table 10).
+  • Rule 2: Conditional simulation of 10,000 observations, depending on the value of gender in each of the 10,000 registries of the churn data, from a Bernoulli with parameter $\mathbb{P}(\text{{exited}}=1\,|\,\text{{gender}}=1)=0{.}1648$ or $\mathbb{P}(\text{{exited}}=1\,|\,\text{{gender}}=0)=0{.}2508,$ see cases 2 and 3 in Table 10.
+  • Rule 3: Conditional simulation of 10,000 observations, depending on the value of active in each of the 10,000 registries of the churn data, from a Bernoulli with parameter according with cases 4 and 5 in Table 10.
+  • Rule 4: Conditional simulation of 10,000 observations, depending on the value of gender and active in each of the 10,000 registries of the churn data, from a Bernoulli with parameter accordingly with cases 6, 7, 8, and 9, in Table 10.
+  • Rule 5: Similar to Rule 4 but fitting a trivariate MBD such that the bivariate dependencies are the same as in Table 9, but setting the trivariate dependence parameter $\mu_{123}=0$ instead of 0.1252.-0{.}1252.
+The purpose of Rule 5 is to compare its accuracy versus Rule 4 and notice the effect of ignoring dependencies beyond pairwise ones, in this case $\mu_{123}.$ For testing Rule 5 we need a vector of 8 trivariate probabilities $\{p_{ijk}:i,j,k\in{0,1}\}$ such that the marginals are the same (that is, same $\theta_{1}=0{.}4543,$ $\theta_{2}=0{.}4851,$ and $\theta_{3}=0{.}7961$ values as in Table 8), the same bivariate dependencies (that is, same $\mu_{12}=0{.}0240,$ $\mu_{13}=-0{.}1916,$ and $\mu_{23}=-0{.}2994$ values as in Table 9), but such that there is no trivariate dependence, that is $\mu_{123}=0.$
+From (27) we notice that $\mu_{123}=0$ if and only if $\theta_{123}=\theta_{1}\theta_{2}\theta_{3}.$ Since $\mu_{12}=0.0240>0$ but $\mu_{13}=-0{.}1916<0$ and $\mu_{23}=-0{.}2994$ then from (9) we may obtain:
+$\theta_{12}$ $=$ $(\min\{\theta_{1},\theta_{2}\}-\theta_{1}\theta_{2})\mu_{12}\,+\,\theta_{1}\theta_{2}\,$,
+$\theta_{13}$ $=$ $(\theta_{1}\theta_{3}-\max\{\theta_{1}+\theta_{3}+1,0\})\mu_{13}\,+\,\theta_{1}\theta_{3}\,$,
+$\theta_{23}$ $=$ $(\theta_{2}\theta_{3}-\max\{\theta_{2}+\theta_{3}+1,0\})\mu_{23}\,+\,\theta_{2}\theta_{3}\,.$
+With the above new values $\theta_{\!A}$ we may apply formulas (12)-(19) to obtain the required probabilities for Rule 5.
+Each of the 5 rules was applied 1 million times to each of the 10,000 registries of the churn data, and for each repetition the percentage of correct predictions compared to the real observed value of exited was calculated, and finally then averaged.
+The results are summarized in Table 11, from which the following remarks are derived:
+  • Since there are significant bivariate dependencies $\mu_{13}\equiv\mu(\text{{gender}},\text{{exited}})=-0{.}1916$ and $\mu_{23}\equiv\mu(\text{{active}},\text{{exited}})=-0{.}2994,$ see Table 9, it was expected an increase of accuracy in predicting exited by just conditioning in one of the variables gender (Rule 2) or active (Rule 3) with respect to not conditioning at all (Rule 1), and the accuracy of Rule 3 is larger than Rule 2 since $|\mu_{23}|>|\mu_{13}|.$
+  • Since both variables gender and active have some predictive power for exited, according to the associated bivariate dependencies $\mu_{13}$ and $\mu_{23},$ it was expected to have even more accuracy by conditioning in both variables simultaneously, so there is no surprise that Rules 4 and 5 have better accuracies than Rules 2 and 3.
+  • Rule 4 has better accuracy than Rule 5, as expected, because the latter ignores the trivariate dependence $\mu_{123}=-0{.}1252.$
+case condition $\mathbb{P}(\text{{exited}}=1\,|\,\text{condition})$
+1 none 0.2039
+2 gender = 1 0.1648
+3 gender = 0 0.2508
+4 active = 1 0.1428
+5 active = 0 0.2686
+6 gender = 1 and active = 1 0.1120
+7 gender = 1 and active = 0 0.2231
+8 gender = 0 and active = 1 0.1814
+9 gender = 0 and active = 0 0.3208
+Table 10: Probabilities for the main variable of interest exited under different conditions, in the trivariate MBD model for the binary variables (gender,active,exited) from bank churn data in [21].
+The Julia code to generate the results is in Appendix A.3.
+Rule # Accuracy
+1 67.54%
+2 67.91%
+3 68.33%
+4 68.69%
+5 68.56%
+Table 11: Accuracy of predictions for exited according to the 5 defined rules in the trivariate MBD model for the binary variables (gender,active,exited) from bank churn data in [21].
+The Julia code to generate the results is in Appendix A.3.
+6.4 A real data example: COVID-19 in Mexico
+Data from official statistics about COVID-19 in Mexico during the first year of the pandemic, available for download from [18] as a covid2020.csv file, with 2.15 million patient registries and 15 binary variables measured (among many others available), with the following description of the variables:
+  1. 1.
+    SEXO = 1 for male, 0 for female.
+  2. 2.
+    TIPO_PACIENTE = 1 for hospitalized, 0 non-hospitalized.
+  3. 3.
+    DIABETES = 1 for diabetic, 0 if not.
+  4. 4.
+    EPOC = 1 for chronic obstructive pulmonary disease (COPD), 0 if not.
+  5. 5.
+    INMUSUPR = 1 for immunosuppression, 0 if not.
+  6. 6.
+    HIPERTENSION = 1 for hypertension, 0 if not.
+  7. 7.
+    CARDIOVASCULAR = 1 for cardiovascular disease, 0 if not.
+  8. 8.
+    OBESIDAD = 1 for obesity, 0 if not.
+  9. 9.
+    RENAL_CRONICA = 1 for chronic kidney disease (CKD), 0 if not.
+  10. 10.
+    TABAQUISMO = 1 for smoker, 0 for non-smoker.
+  11. 11.
+    e00 = 1 for ages between 0 and 19, 0 if not.
+  12. 12.
+    e20 = 1 for ages between 20 and 39, 0 if not.
+  13. 13.
+    e40 = 1 for ages between 40 and 64, 0 if not.
+  14. 14.
+    e65 = 1 for ages between 65 and older, 0 if not.
+  15. 15.
+    MUERTE = 1 for death, 0 if survived.
+In this case $n=15$ and therefore there are $2^{n}-1=32,768$ probabilities to estimate, $2^{n}-1$ parameters $\theta_{A}$ to estimate, and $2^{n}-(n+1)$ multivariate dependencies $\mu_{A}$ to estimate (pairwise, three-wise, up to 15-wise).
+The Julia code to fit a MBD is in Appendix A.4.
+In Table 12 it is calculated the 20 largest values for dependence parameters that include variable 15 (death), from which it is possible to identify in all cases the presence of variable 2 (hospitalization), and in several cases variables 4, 9, 5, 7 and 14 (COPD, CKD, immunosuppression, cardiovascular disease, and age 65 or older).
+point estimate
+$A\subseteq\{1,\ldots,15\}$ $\mu_{A}$
+$\{2,15\}$ 0.9021
+$\{2,4,15\}$ 0.8590
+$\{2,9,15\}$ 0.8581
+$\{2,5,15\}$ 0.8550
+$\{2,4,9,15\}$ 0.8311
+$\{2,7,15\}$ 0.8310
+$\{2,5,9,15\}$ 0.8283
+$\{2,4,5,15\}$ 0.8272
+$\{2,7,9,15\}$ 0.8094
+$\{2,4,5,9,15\}$ 0.8090
+$\{2,4,7,15\}$ 0.8088
+$\{2,5,7,15\}$ 0.8032
+$\{2,4,7,9,15\}$ 0.7940
+$\{2,5,7,9,15\}$ 0.7897
+$\{2,4,5,7,15\}$ 0.7879
+$\{2,4,5,7,9,15\}$ 0.7784
+$\{2,4,9,14,15\}$ 0.6594
+$\{2,4,14,15\}$ 0.6583
+$\{2,14,15\}$ 0.6565
+$\{2,9,14,15\}$ 0.6560
+Table 12: The 20 largest values for dependence parameters that include variable 15 (death). For details of the calculations see Appendix A.4.
+It was of main interest to identify the variables more closely related to the fatal outcome of death (variable 15).
+This is specially relevant because conditioning in such variables a better assessment of mortality risk could be obtained.
+In this data the (non-conditional) mortality rate was $10.1\%$ but conditioning on variables 2 (hospitalization), 4 (COPD), 9 (CKD), 5 (immunosupression), 7 (cardiovascular) and 14 (age 65 and older), it increases up to $66.3\%$ (for details of the calculations see Appendix A.4).
+7 Conclusions
+In this paper, a novel subcopula-based characterization of dependence for the MBD has been introduced.
+By leveraging Sklar’s theorem, it has been shown that subcopulas can effectively capture the dependence structure between Bernoulli random variables, extending the utility of subcopula theory to binary data.
+The proposed approach provides a flexible framework for studying multivariate binary data by separating the marginal distributions from the joint dependence structure.
+Explicit formulas and bounds for both bivariate and multivariate Bernoulli distributions in terms of subcopulas were derived, allowing for a more granular understanding of the dependence parameters.
+In particular, dependence measures were obtained for bivariate and trivariate Bernoulli distributions, and proposed a general method to extend these results to higher dimensions.
+This work’s findings confirm that subcopulas are a powerful tool for modeling and quantifying dependence of all orders in the MBD, not just pairwise.
+The proposed methodology not only advances the theoretical understanding of subcopulas but also opens up new possibilities for practical applications in fields dealing with multivariate binary data.
+For example, in epidemiology, the framework can model the joint occurrence of comorbidities in infectious disease outbreaks, such as diabetes, hypertension, COPD, and mortality in COVID-19 patients, or assess the co-occurrence of vaccine side effects.
+In genetics, it can be applied to analyze dependencies among binary outcomes such as gene activation/inactivation, the presence or absence of mutations, or epistatic interactions between SNPs (Single Nucleotide Polymorphisms).
+In finance, the approach is directly relevant for credit risk modeling, where default/no-default indicators across obligors must be analyzed jointly, as well as for systemic risk assessment in portfolios, where simultaneous threshold-crossing events in asset returns are of interest.
+Moreover, the Bayesian inference framework presented here offers robust tools for parameter estimation and uncertainty quantification, enhancing the practical implementation of the proposed approach.
+Computational implementations of the MBD are quite scarce, so the set of functions provided in Section 6 and downloadable from [18] represents a contribution that facilitates the analysis and simulation of multivariate binary data.
+Future research could explore further generalizations of the subcopula approach to other types of discrete data and investigate its performance in applied contexts.
+Additionally, incorporating covariates or other forms of heterogeneity into the subcopula framework may provide more nuanced models, particularly for data sets with more complex dependence structures.
+Appendix A Julia code for the examples
+All the source code is available in a public repository at [18] for full reproducibility, and must be loaded and executed following the provided indications.
+A.1 A theoretical example (6.1)
+Ψ
+Ψ# setup parameters and MBD model
+Ψpp = [0.15,0.21,0.21,0.03,0.21,0.03,0.03,0.13];
+ΨX = MBerDep(pp);
+Ψ
+Ψ# parameters
+Ψ[X.dparam.idx X.dparam.value]
+Ψ
+Ψ# dependence measures
+Ψ[X.dmeas.idx X.dmeas.value]
+Ψ
+A.2 A simulation example (6.2)
+Define a trivariate Bernoulli distribution applying Algorithm 1 from Section 4:
+    begin
     p000 = 0.1
@@
     pp = [p000, p001, p010, p011, p100, p101, p110, p111]
+    X = MBerDep(pp)
     end
-8-element Vector{Float64}:
- 0.1
- 0.2
- 0.1
- 0.2
- 0.05
- 0.15
- 0.1
- 0.09999999999999987
-julia> X = MBerDep(pp); X.nrvs
-3
-where X.nrvs is the corresponding dimension of the random vector.
-Then we simulate a random sample of size $3,000$ and make point and interval estimations about the probability parameters, and the dependence parameters and measures, and a comparison to the theoretical values.
-In the estimation matrices the third column is always a median posterior point estimation, and the second and fourth columns are the extreme points of a $99\%$ probability posterior interval estimation:
-julia> begin
+Simulate a random sample of size $3,000$:
+    begin
     Random.seed!(1234) # for reproducibility
     simX = MBerSim(pp, 3_000);
+    end
+Interval and point estimations for the probability parameters:
     infer = MBerInf(simX, prior = 1/2, nsim = 10_000, probint = 0.99);
-           infer.probs # estimated probabilities
+    infer.probs # estimations
+    [X.binprob.idx X.binprob.value] # theoretical probabilities
+Interval and point estimations for the dependence parameters:
+    begin
+    infer.dparam # estimations
+    [X.dparam.idx X.dparam.value] # theoretical dependence parameters
     end
-84 Matrix{Any}:
- [0, 0, 0]  0.0861693  0.0996407  0.1143
- [0, 0, 1]  0.176657   0.19481    0.21376
- [0, 1, 0]  0.087351   0.100971   0.115828
- [0, 1, 1]  0.175131   0.193508   0.212438
- [1, 0, 0]  0.0332592  0.0420294  0.0520198
- [1, 0, 1]  0.139136   0.155535   0.173265
- [1, 1, 0]  0.090522   0.104274   0.119274
- [1, 1, 1]  0.0945739  0.108592   0.1238
-julia> [X.binprob.idx X.binprob.value] # theoretical probabilities
-82 Matrix{Any}:
- [0, 0, 0]  0.1
- [0, 0, 1]  0.2
- [0, 1, 0]  0.1
- [0, 1, 1]  0.2
- [1, 0, 0]  0.05
- [1, 0, 1]  0.15
- [1, 1, 0]  0.1
- [1, 1, 1]  0.1
-julia> infer.dparam # estimated dependence parameters
-74 Matrix{Any}:
- [1, 2, 3]  0.0861693  0.0996407  0.1143
- [1, 2]     0.273408   0.294522   0.316158
- [1, 3]     0.182337   0.200692   0.219919
- [2, 3]     0.125958   0.141746   0.158641
- [1]        0.566172   0.589215   0.612288
- [2]        0.468764   0.492305   0.515974
- [3]        0.325026   0.347202   0.369748
-julia> [X.dparam.idx X.dparam.value] # theoretical dependence parameters
-72 Matrix{Any}:
- [1, 2, 3]  0.1
- [1, 2]     0.3
- [1, 3]     0.2
- [2, 3]     0.15
- [1]        0.6
- [2]        0.5
- [3]        0.35
-julia> infer.dmeas # estimated dependence measures
-44 Matrix{Any}:
- [1, 2, 3]  -0.113144   -0.0104273   0.039112
- [1, 2]     -0.0338129   0.0222768   0.0793568
- [1, 3]     -0.072454   -0.0188205   0.0504544
- [2, 3]     -0.23491    -0.170452   -0.105142
-julia> [X.dmeas.idx X.dmeas.value] # theoretical dependence measures
-42 Matrix{Any}:
- [1, 2, 3]  -0.047619
- [1, 2]      0.0
- [1, 3]     -0.047619
- [2, 3]     -0.142857
-Example 4.
-Data from official statistics about COVID-19 in Mexico during the first year of the pandemic, available for download from [4] as a covid2020.csv file, with 2.15 million patient registries and 15 binary variables measured (among many others available).
+Coverage rates for all parameters:
+    @time begin # WARNING: 10_000 simulations take around 3.5 hours
+      nsim = 10_000
+      nobs = 3_000
+      cover_probs = zeros(Bool, nsim, 8)
+      cover_param = zeros(Bool, nsim, 7)
+      cover_meas = zeros(Bool, nsim, 4)
+      for i in 1:nsim
+         simX = MBerSim(pp, nobs)
+         infer = MBerInf(simX,prior=1/2,nsim=100000,probint=0.99)
+         for j in 1:8
+           cover_probs[i,j]=(infer.probs[j,2]<=pp[j]<=infer.probs[j,4])
+         end
+         for j in 1:7
+           cover_param[i,j]=(infer.dparam[j,2]<=X.dparam.value[j]
+           <=infer.dparam[j,4])
+         end
+         for j in 1:4
+           cover_meas[i,j]=(infer.dmeas[j,2]<=X.dmeas.value[j]
+           <=infer.dmeas[j,4])
+         end
+      end
+      println("Coverage rates (99% intervals):")
+      println("Probabilities: ", mean(cover_probs, dims = 1))
+      println("Parameters:    ", mean(cover_param, dims = 1))
+      println("Dependencies:  ", mean(cover_meas, dims = 1))
+      println()
+    end
+A.3 A real data example: Bank churn data (6.3)
+    begin # Read data as a dataframe
+        println("Example Section 6.4: Bank churn data")
+        df = CSV.read("churnbindata.csv", DataFrame)
+        show(describe(df), allrows = true)
+    end
+    begin # Convert dataframe to a matrix
+        binmat = zeros(Int, size(df))
+        for c in 1:ncol(df)
+            binmat[:, c] = df[:, c]
+        end
+        binmat
+    end
+    # Point and 95% probability interval estimations
+    @time inference4v = MBerInf(binmat, nsim = 1_000_000)
+    inference4v.dmeas
+    # Discard variable ’crcard’ (column 2)
+    binmat = binmat[:, [1,3,4]]
+    @time inference3v = MBerInf(binmat, nsim = 1_000_000)
+    # Point and 95% probability interval estimations for...
+    inference3v.probs # ... probabilities
+    inference3v.dparam # ... parameters
+    inference3v.dmeas # ... dependencies
+    # Probabilities for exiting the bank
     begin
-    df = CSV.read("covid2020.csv", DataFrame)
-    show(describe(df), allrows = true)
-    data = zeros(Int, size(df))
-    for c in 1:ncol(df)
-        data[:, c] = df[:, c]
+    condProb = DataFrame(gender = [missing,1,0,missing,missing,1,1,0,0],
+                         active = [missing,missing,missing,1,0,1,0,1,0],
+                         Pexit = zeros(9))
+    condProb.Pexit[1] = 1 - inference3v.dparam[7,3]
+    # warning due to rounding
+    X123 = MBerDep(Float64.(inference3v.probs[:, 3]))
+    condProb.Pexit[2] = MBerCond(X123.binprob.value, [3], [1],
+    [1]).binprob.value[2]
+    condProb.Pexit[3] = MBerCond(X123.binprob.value, [3], [1],
+    [0]).binprob.value[2]
+    condProb.Pexit[4] = MBerCond(X123.binprob.value, [3], [2],
+    [1]).binprob.value[2]
+    condProb.Pexit[5] = MBerCond(X123.binprob.value, [3], [2],
+    [0]).binprob.value[2]
+    condProb.Pexit[6] = MBerCond(X123.binprob.value, [3], [1,2],
+    [1,1]).binprob.value[2]
+    condProb.Pexit[7] = MBerCond(X123.binprob.value, [3], [1,2],
+    [1,0]).binprob.value[2]
+    condProb.Pexit[8] = MBerCond(X123.binprob.value, [3], [1,2],
+    [0,1]).binprob.value[2]
+    condProb.Pexit[9] = MBerCond(X123.binprob.value, [3], [1,2],
+    [0,0]).binprob.value[2]
+    display(condProb)
     end
-    println()
-    data
+# Prediction rules
+function predNonCond(nsim, condProb)
+    B = Bernoulli(condProb.Pexit[1])
+    accuracy = 0.0
+    for k in 1:nsim
+        accuracy += mean(rand(B, 10_000) .== binmat[:, 3]) / nsim
     end
-15x7 DataFrame
- Row | variable        mean        min    median   max    nmissing  eltype
-     | Symbol          Float64     Int64  Float64  Int64  Int64     DataType
-   1 | SEXO            0.501803        0      1.0      1         0  Int64
-   2 | TIPO_PACIENTE   0.196705        0      0.0      1         0  Int64
-   3 | DIABETES        0.1348          0      0.0      1         0  Int64
-   4 | EPOC            0.0113235       0      0.0      1         0  Int64
-   5 | INMUSUPR        0.00843405      0      0.0      1         0  Int64
-   6 | HIPERTENSION    0.174666        0      0.0      1         0  Int64
-   7 | CARDIOVASCULAR  0.0158313       0      0.0      1         0  Int64
-   8 | OBESIDAD        0.144864        0      0.0      1         0  Int64
-   9 | RENAL_CRONICA   0.0152094       0      0.0      1         0  Int64
-  10 | TABAQUISMO      0.0735159       0      0.0      1         0  Int64
-  11 | e00             0.0502618       0      0.0      1         0  Int64
-  12 | e20             0.38682         0      0.0      1         0  Int64
-  13 | e40             0.434831        0      0.0      1         0  Int64
-  14 | e65             0.128087        0      0.0      1         0  Int64
-  15 | MUERTE          0.101051        0      0.0      1         0  Int64
-2154719x15 Matrix{Int64}:
- 0  0  1  0  0  0  0  0  0  0  0  0  1  0  0
- 1  0  0  0  0  0  0  0  0  0  0  1  0  0  0
- 0  0  1  0  0  0  0  0  0  0  0  0  1  0  0
- 1  1  1  1  0  0  1  1  0  0  0  0  0  1  1
- 0  1  1  0  0  1  0  1  0  0  0  0  1  0  1
- 1  1  0  0  0  0  0  1  0  0  0  0  0  1  1
-With the following description of the variables:
-  1. 1.
-    SEXO = $1$ for male, 00 for female.
-  2. 2.
-    TIPO_PACIENTE = $1$ for hospitalized, 00 non-hospitalized.
-  3. 3.
-    DIABETES = $1$ for diabetic, 00 if not.
-  4. 4.
-    EPOC = $1$ for chronic obstructive pulmonary disease (COPD), 00 if not.
-  5. 5.
-    INMUSUPR = $1$ for immunosuppression, 00 if not.
-  6. 6.
-    HIPERTENSION = $1$ for hypertension, 00 if not.
-  7. 7.
-    CARDIOVASCULAR = $1$ for cardiovascular disease, 00 if not.
-  8. 8.
-    OBESIDAD = $1$ for obesity, 00 if not.
-  9. 9.
-    RENAL_CRONICA = $1$ for chronic kidney disease (CKD), 00 if not.
-  10. 10.
-    TABAQUISMO = $1$ for smoker, 00 for non-smoker.
-  11. 11.
-    e00 = $1$ for ages between 00 and $19$, 00 if not.
-  12. 12.
-    e20 = $1$ for ages between $20$ and $39$, 00 if not.
-  13. 13.
-    e40 = $1$ for ages between $40$ and $64$, 00 if not.
-  14. 14.
-    e65 = $1$ for ages between $65$ and older, 00 if not.
-  15. 15.
-    MUERTE = $1$ for death, 00 if survived.
-In this case $n=15$ and therefore there are $7672^{15}-1=32,$ 767 probabilities to estimate, $12^{15}-$ 1 parameters to estimate, and $162^{15}-$ 16 multivariate dependencies to estimate (pairwise, three-wise, up to 15-wise).
-@time estim = MBerEst(data); # 5 minutes approx
-If we are interested in the 20 largest values for dependence parameters that include variable 15 (death) then:
+    return accuracy
+end
+function predGivenGender(nsim, condProb)
+    ival1 = findall(binmat[:, 1] .== 1)
+    ival0 = findall(binmat[:, 1] .== 0)
+    n1 = length(ival1)
+    n0 = length(ival0)
+    vsim = fill(-99999999, 10_000)
+    accuracy = 0.0
+    for k in 1:nsim
+        vsim[ival1] = rand(Bernoulli(condProb.Pexit[2]), n1)
+        vsim[ival0] = rand(Bernoulli(condProb.Pexit[3]), n0)
+        accuracy += mean(vsim .== binmat[:, 3]) / nsim
+    end
+    return accuracy
+end
+function predGivenActive(nsim, condProb)
+    ival1 = findall(binmat[:, 2] .== 1)
+    ival0 = findall(binmat[:, 2] .== 0)
+    n1 = length(ival1)
+    n0 = length(ival0)
+    vsim = fill(-99999999, 10_000)
+    accuracy = 0.0
+    for k in 1:nsim
+        vsim[ival1] = rand(Bernoulli(condProb.Pexit[4]), n1)
+        vsim[ival0] = rand(Bernoulli(condProb.Pexit[5]), n0)
+        accuracy += mean(vsim .== binmat[:, 3]) / nsim
+    end
+    return accuracy
+end
+function predGivenGenderActive(nsim, condProb)
+    X1X2 = Vector{Int}[]
+    for j in 1:10_000
+        push!(X1X2, binmat[j, [1,2]])
+    end
+    ival11 = findall(X1X2 .== [[1,1]])
+    ival10 = findall(X1X2 .== [[1,0]])
+    ival01 = findall(X1X2 .== [[0,1]])
+    ival00 = findall(X1X2 .== [[0,0]])
+    n11 = length(ival11)
+    n10 = length(ival10)
+    n01 = length(ival01)
+    n00 = length(ival00)
+    vsim = fill(-99999999, 10_000)
+    accuracy = 0.0
+    for k in 1:nsim
+        vsim[ival11] = rand(Bernoulli(condProb.Pexit[6]), n11)
+        vsim[ival10] = rand(Bernoulli(condProb.Pexit[7]), n10)
+        vsim[ival01] = rand(Bernoulli(condProb.Pexit[8]), n01)
+        vsim[ival00] = rand(Bernoulli(condProb.Pexit[9]), n00)
+        accuracy += mean(vsim .== binmat[:, 3]) / nsim
+    end
+    return accuracy
+end
+# Accuracy for each rule
+nsim = 1_000_000 # 4 minutes approx
+@time predict=DataFrame(rule=["Non-conditional","Given gender",
+                      "Given active", "Given gender & active"],
+            accuracy = [predNonCond(nsim, condProb),
+                        predGivenGender(nsim, condProb),
+                        predGivenActive(nsim, condProb),
+                        predGivenGenderActive(nsim, condProb)])
+# Accuracy given gender and active status
+# but setting trivariate dependence to zero
+begin
+    th1 = inference3v.dparam[5,3]
+    th2 = inference3v.dparam[6,3]
+    th3 = inference3v.dparam[7,3]
+    th123 = th1 * th2 * th3
+    mu12 = inference3v.dmeas[2,3]
+    mu13 = inference3v.dmeas[3,3]
+    mu23 = inference3v.dmeas[4,3]
+    th12 = (min(th1,th2) - th1*th2)*mu12 + th1*th2
+    th13 = (th1*th3 - max(th1+th3-1,0))*mu13 + th1*th3
+    th23 = (th2*th3 - max(th2+th3-1,0))*mu23 + th2*th3
+    p000 = th123
+    p001 = th12 - th123
+    p010 = th13 - th123
+    p100 = th23 - th123
+    p011 = th1 - th12 - th13 + th123
+    p101 = th2 - th12 - th23 + th123
+    p110 = th3 - th13 - th23 + th123
+    p111 = 1 - (p000 + p001 + p010 + p011 + p100 + p101 + p110)
+    pp = [p000, p001, p010, p011, p100, p101, p110, p111]
+    Xzero3dep = MBerDep(pp)
+end;
+# dependence measures without and with zero trivariate dependence
+[inference3v.dmeas[:, [1,3]] Xzero3dep.dmeas.value]
+# Probabilities for exiting the bank with zero trivariate dependence
+begin
+condProb3 = DataFrame(gender = [missing,1,0,missing,missing,1,1,0,0],
+                      active = [missing,missing,missing,1,0,1,0,1,0],
+                      Pexit = zeros(9))
+condProb3.Pexit[1] = 1 - Xzero3dep.dparam.value[7]
+condProb3.Pexit[2] = MBerCond(Xzero3dep.binprob.value, [3], [1],
+[1]).binprob.value[2]
+condProb3.Pexit[3] = MBerCond(Xzero3dep.binprob.value, [3], [1],
+[0]).binprob.value[2]
+condProb3.Pexit[4] = MBerCond(Xzero3dep.binprob.value, [3], [2],
+[1]).binprob.value[2]
+condProb3.Pexit[5] = MBerCond(Xzero3dep.binprob.value, [3], [2],
+[0]).binprob.value[2]
+condProb3.Pexit[6] = MBerCond(Xzero3dep.binprob.value, [3], [1,2],
+[1,1]).binprob.value[2]
+condProb3.Pexit[7] = MBerCond(Xzero3dep.binprob.value, [3], [1,2],
+[1,0]).binprob.value[2]
+condProb3.Pexit[8] = MBerCond(Xzero3dep.binprob.value, [3], [1,2],
+[0,1]).binprob.value[2]
+condProb3.Pexit[9] = MBerCond(Xzero3dep.binprob.value, [3], [1,2],
+[0,0]).binprob.value[2]
+display(condProb3)
+end
+# Accuracy for each rule setting trivariate dependence to zero
+nsim = 1_000_000 # 4 minutes approx
+@time predict3=DataFrame(rule=["Non-conditional","Given gender",
+                         "Given active","Given gender & active"],
+                accuracy = [predNonCond(nsim, condProb3),
+                        predGivenGender(nsim, condProb3),
+                        predGivenActive(nsim, condProb3),
+                        predGivenGenderActive(nsim, condProb3)]);
+# Comparison of accuracies with and without trivariate dependence
+println("Accuracy with trivariate dependence:")
+predict
+println("Accuracy with zero trivariate dependence:")
+predict3
+A.4 A real data example: COVID-19 in Mexico (6.4)
+Ψ
+Ψbegin # load data
+ΨΨdf = CSV.read("covid2020.csv", DataFrame)
+ΨΨshow(describe(df), allrows = true)
+ΨΨdata = zeros(Int, size(df))
+ΨΨfor c in 1:ncol(df)
+ΨΨΨdata[:, c] = df[:, c]
+ΨΨend
+ΨΨprintln()
+ΨΨdata
+Ψend
+Ψ
+@time estim = MBerEst(data); # Fit MBD model (5 minutes approx)Ψ
+# 20 largest values for dependence parameters
+# that include variable 15 (death)
 begin
@@
 m = estim.dmeas.value[iord]
-    iMue = findall(x -> 15 in x, midx)
-    [midx[iMue] m[iMue]][1:20, :]
+iMue = findall(x -> 15 in x, midx)[midx[iMue] m[iMue]][1:20, :]
 end
-202 Matrix{Any}:
- [2, 15]              0.902144
- [2, 4, 15]           0.859008
- [2, 9, 15]           0.858144
- [2, 5, 15]           0.855051
- [2, 4, 9, 15]        0.831122
- [2, 7, 15]           0.831098
- [2, 5, 9, 15]        0.828303
- [2, 4, 5, 15]        0.827159
- [2, 7, 9, 15]        0.809365
- [2, 4, 5, 9, 15]     0.808968
- [2, 4, 7, 15]        0.808805
- [2, 5, 7, 15]        0.803182
- [2, 4, 7, 9, 15]     0.793962
- [2, 5, 7, 9, 15]     0.789702
- [2, 4, 5, 7, 15]     0.787861
- [2, 4, 5, 7, 9, 15]  0.778416
- [2, 4, 9, 14, 15]    0.659362
- [2, 4, 14, 15]       0.658259
- [2, 14, 15]          0.656512
- [2, 9, 14, 15]       0.65601
-Therefore from the 20 highest values for multivariate dependence that always include death we are able to identify in all cases variable 2 (hospitalization), and in several cases variables 4, 9, 5, 7 and 14 (COPD, CKD, immunosuppression, cardiovascular disease, and age 65 or older).
-In this data the mortality rate was $10.1\%$ but conditioning on the mentioned variables it increases up to $66.3\%:$
-begin
-    pd = mean(data[:, 15])
+Ψ
+begin # conditional risk
+    pd = mean(data[:, 15]) # general mortality rate
     estim2 = MBerEst(data[:, [2,4,5,7,9,14,15]])
-    pdcond = MBerCond(estim2.binprob.value, [7], [1,2,3,4,5,6],
-                      [1,1,1,1,1,1]).binprob.dic[[1]]
+    # conditional mortality rate
+    pdcond = MBerCond(estim2.binprob.value, [7], [1,2,3,4,5,6],Ψ
+Ψ       [1,1,1,1,1,1]).binprob.dic[[1]]
     println("P(death) = ", pd)
-    println("P(death|hospital, COPD, immunosup, cardio, CKD,
-             age 65+) = ", pdcond)
+    println("P(death|hospital, COPD, immunosup, cardio,
+Ψ      CKD, age 65+) = ", pdcond)
 end
-P(death) = 0.10105122756145929
-P(death|hosp,COPD,immunosup,cardio,CKD,65+) = 0.6632653061224489
-6 Conclusions
-In this paper, a novel subcopula-based characterization of dependence for the Multivariate Bernoulli Distribution (MBD) has been introduced.
-By leveraging Sklar’s theorem, it has been shown that subcopulas can effectively capture the dependence structure between Bernoulli random variables, extending the utility of subcopula theory to binary data.
-The proposed approach provides a flexible framework for studying multivariate binary data by separating the marginal distributions from the joint dependence structure.
-Explicit formulas and bounds for both bivariate and multivariate Bernoulli distributions in terms of subcopulas were derived, allowing for a more granular understanding of the dependence parameters.
-In particular, dependence measures were obtained for bivariate and trivariate Bernoulli distributions, and proposed a general method to extend these results to higher dimensions.
-This work’s findings confirm that subcopulas are a powerful tool for modeling and quantifying dependence in the MBD of all orders, not just pairwise.
-The proposed methodology not only advances the theoretical understanding of subcopulas but also opens up new possibilities for practical applications in fields dealing with multivariate binary data.
-Potential applications include genetics, epidemiology, and social sciences, where understanding the dependence structure between binary variables is essential.
-Moreover, the bayesian inference framework presented offers a robust tool for parameter estimation and uncertainty quantification, enhancing the practical implementation of the proposed approach.
-Computational implementations of the MBD are quite scarce, so the set of functions provided in subsection 5.2 and downloadable from [4] is a contribution that may help in analyzing and simulating multivariate binary data.
-Future research could explore further generalizations of the subcopula approach to other types of discrete data and investigate its performance in applied contexts.
-Additionally, incorporating covariates or other forms of heterogeneity into the subcopula framework may provide more nuanced models, particularly for data sets with more complex dependence structures.
+DECLARATIONS
+Ethics approval and consent to participate
+Not applicable. This study did not involve human participants, animals, or sensitive personal data requiring ethical approval.
+Consent for publication
+Not applicable. This manuscript does not contain data from any individual person.
+Availability of data and materials
+All the data and programming code is openly available at the following Github repository where the author of this repository is the same as the submitted contribution: https://github.com/aerdely/MultivariateBernoulli
+Competing interests
+No, I declare that the authors have no competing interests as defined by Springer, or other interests that might be perceived to influence the results and/or discussion reported in this paper.
+Funding
+Dirección General de Asuntos del Personal Académico, Universidad Nacional Autónoma de México, PAPIIT IN104425.
+Authors’ contributions
+As a single author A.E has worked all the details in this contribution.
+Acknowledgements
+The author gratefully acknowledges the reviewer whose comments helped to significantly improve this work.